### v1

# Renormalon-based resummation for spacelike and timelike QCD quantities whose perturbation expansion has general form

Abstract
We present a generalisation of our previous approach of a renormalon-motivated resummation of the QCD observables. Previously it was applied to the spacelike observables whose perturbation expansion was ${\cal D}(Q^{2})=a(Q^{2})+{\cal O}(a^{2})$, where $a(Q^{2})\equiv\alpha_{s}(Q^{2})/\pi$ is the running QCD coupling. Now we generalise the resummation to spacelike quantities ${\cal D}(Q^{2})=a(Q^{2})^{\nu_{0}}+{\cal O}(a^{\nu_{0}+1})$ and timelike quantities ${\cal F}(\sigma)=a(\sigma)^{\nu_{0}}+{\cal O}(a^{\nu_{0}+1})$, where $\nu_{0}$ is in general a noninteger number ($0<\nu_{0}\leq 1$). We evaluate with this approach a timelike quantity, namely the scheme-invariant factor of the Wilson coefficient of the chromomagnetic operator in the heavy-quark effective Lagrangian, and related quantities.
renormalons; resummations; perturbative QCD; QCD phenomenology; holomorphic QCD
I Introduction
The study of QCD properties at low-momentum transfers ($\lesssim 1$ GeV) is challenging due to the nonperturbative nature of QCD. There are some efforts to try to unify the high- and low-momentum regimes, and most of them lead to noncontinuous and/or nonoholomorphic transtion between these two regimes. Recently, in BrodECH ; ACECH , certain effective charges with holomorphic connection between the two regimes are supposed to contain information of the theory, but they are unfortunately tied to only one specific observable.
In this work, we present a resummation formalism that evaluates the (leading-twist part of the) QCD observables with a single integral, keeping some of the ideas of the effective charge, but with clear differences. An early version of this resummation was constructed by Neubert Neubert some time ago for the large-$\beta_{0}$ approximation (cf. also NC ). Subsequently, the resummation formalism was extended to any loop-level renmod , leading to an integration involving the (exact) running coupling and an observable-dependent characteristic (weight) function, where the latter can be (approximately) determined by the knowledge of the renormalon structure of the considered observable. These resummations were constructed for the case when the perturbation expansions of the considered observables were series of integer powers of the coupling. In the present work, we extend these resummations to the case of observables whose perturbation expansions have, in general, noninteger powers of the couplings. In principle, such resummations are for the spacelike QCD observables, but here we also extend this formalism to the general case of timelike QCD observables, and present a phenomenological application of this formalism.
In Ref. renmod , a renormalon-motivated resummation procedure has been developed for spacelike QCD observables whose perturbation expansion is ${\cal D}(Q^{2})=a(Q^{2})+{\cal O}(a^{2})$, where $a(Q^{2})\equiv\alpha_{s}(Q^{2})/\pi$ and $Q^{2}\equiv-q^{2}$ ($(q^{0})^{2}+{\vec{q}}^{2}$) is in the spacelike (i.e., non-timelike) regime in the complex plane, i.e., $Q^{2}\in\mathbb{C}\backslash(-\infty,0)$. Furthermore, a simple extension to timelike variables ${\cal F}(\sigma)$ ($\sigma>0$) was made there, for the case when ${\cal F}(\sigma)$ is a contour integral of the aforementioned spacelike ${\cal D}(Q^{2})$. The expansion of ${\cal D}(Q^{2})$ was considered to be known exactly. In renmod the resummation was applied to the Adler function ${\cal D}(Q^{2})=d(Q^{2})_{\rm Adl}$ and to the semihadronic $\tau$-lepton decay ratio ${\cal F}(\sigma)=r_{\tau}(\sigma)$ ($\sigma=m_{\tau}^{2}$). The resummation method was applied with perturbative (pQCD) coupling $a(Q^{2})$, i.e., a coupling that has Landau singularities in the spacelike region $0\leq Q^{2}\leq{\Lambda}_{c}^{2}$. Thereafter, the same procedure was also performed there with holomorphic QCD couplings ${\mathcal{A}}(Q^{2})$ (i.e., when $a(Q^{2})$ is replaced by ${\mathcal{A}}(Q^{2})$), i.e., couplings that have no Landau singularities and are thus holomorphic for $Q^{2}\in\mathbb{C}\backslash(-\infty,0)$ and hence reflect correctly the holomorphic structure of the spacelike QCD observables ${\cal D}(Q^{2})$ in the complex $Q^{2}$-plane.
Later, we applied this method specifically to the (spacelike) QCD quantity Bjorken polarised sume rule (BSR), $d(Q^{2})_{\rm BSR}$, with the coupling being either pQCD BSRpQCD or holomorphic BSRAQCD . With the holomorphic coupling, we can describe successfully the experimental data for $d(Q^{2})_{\rm BSR}$ in a wider interval of positive $Q^{2}$ that includes also $Q^{2}<1\ {\rm GeV}^{2}$.
In this work, we extend the described method to the resummation of the spacelike observables whose perturbation expansion is ${\cal D}=a(Q^{2})^{\nu_{0}}+{\cal O}(a^{\nu_{0}+1})$ and timelike quantities with perturbation expansion ${\cal F}(\sigma)=a(\sigma)^{\nu_{0}}+{\cal O}(a^{\nu_{0}+1})$, where $\nu_{0}$ has in general a noninteger value ($0<\nu_{0}\leq 1$). As an illustration, we will apply this approach to a specific timelike observable ${\cal F}(\sigma)$ (with $\nu_{0}=1/3$), and will perform the renormalon-motivated resummation by using both the pQCD coupling $a$ and a holomorphic coupling $\mathcal{A}$}.
II Resummation for the case of the spacelike observable ${\cal D}(Q^{2})$
We will describe the formalism by first using the pQCD coupling $a(Q^{2})$, and thereafter comment on the simple changes that have to be made when the holomorphic coupling ${\mathcal{A}}(Q^{2})$ is used.
The perturbation expansion of the considered spacelike obervable ${\cal D}(Q^{2})$ is
${\cal D}(Q^{2})$ $=$ $\sum_{n=0}^{\infty}d_{n}(\nu_{0};\kappa)a(\kappa Q^{2})^{\nu_{0}+n}$ (1a)
$=$ $\sum_{n=0}^{\infty}{\widetilde{d}}_{n}(\nu_{0};\kappa){\widetilde{a}}_{\nu_{0}+n}(\kappa Q^{2}).$ (1b)
Here, $\nu_{0}$ is the index of the first expansion term ($0<\nu_{0}\leq 1$), and $\kappa\equiv\mu^{2}/Q^{2}$ denotes the renormalisation scale parameter which is positive and chosen to be $\kappa\sim 1$. The first expansion is in powers of $a(\kappa Q^{2})$, with the expansion coefficients $d_{n}(\nu_{0};\kappa)$ (note: $d_{0}=1$ by convention). The second expansion is in couplings ${\widetilde{a}}_{\nu_{0}+n}$ which are the noninteger generalisation of the logarithmic derivatives of $a$, with the expansion coefficients ${\widetilde{d}}_{n}(\nu_{0};\kappa)$ (note: ${\widetilde{d}}_{0}=1$).111 For the integer $\nu_{0}=1$ we have: ${\widetilde{a}}_{n+1}(Q^{2})\equiv(-1)^{n}(n!\beta_{0}^{n})^{-1}(d/d\ln Q^{2})^{n}a(Q^{2})$. The noninteger powers and the noninteger logarithmic derivatives are related GCAK
${\widetilde{a}}_{\nu}$ $=$ $\sum_{m=0}^{\infty}k_{m}(\nu)a^{\nu+m}\qquad(k_{0}(\nu)=1)$ (2a)
$a^{\nu}$ $=$ $\sum_{m=0}^{\infty}{{\widetilde{k}}}_{m}(\nu){\widetilde{a}}_{\nu+m}\qquad({\widetilde{k}}_{0}(\nu)=1).$ (2b)
Explicit expressions for the coefficients $k_{m}(\nu)$ and ${{\widetilde{k}}}_{m}(\nu)$, valid for any $\nu>0$ ($\nu$ could be integer) and for $m\leq 4$, are given in GCAK  [^2] In the first line of Eq. (A.11) of App. A in GCAK there is an (obvious) typo, a sign is missing, the correct equation is ${\tilde{k}}_{1}(\nu)=-k_{1}(\nu)$. and are independent of the momentum $Q^{2}$ or $\kappa Q^{2}$ in $a\equiv a(\kappa Q^{2})$. On the other hand, when the coupling is instead holomorphic ($a\mapsto{\mathcal{A}}$), the general explicit expression for the holomorphic analog ${\widetilde{\mathcal{A}}}_{\nu}(Q^{2})$ of ${\widetilde{a}}_{\nu}(Q^{2})$, for any $\nu>-1$ and for any holomorphic QCD (AQCD) framework, was derived in Ref. GCAK (see later). [^3] We point out already here that in AQCD frameworks the couplings $\mathcal{A}$} (${\widetilde{\mathcal{A}}}_{1}$) and ${\widetilde{\mathcal{A}}}_{\nu}$ are the basic couplings, and from them we derive the power analogs ${\mathcal{A}}_{\nu}$ (of $a^{\nu}$). This is just the opposite to the situation in pQCD, where the basic couplings are $a$ and the powers $a^{\nu}$, and from them the couplings ${\widetilde{a}}_{\nu}$ are derived, cf. Eq. (2a).
When we use the relation (2b) in the expansion (1a), and take into account the notation (1b), we obtain the following relations between the coefficients ${\widetilde{d}}_{n}$ and $d_{k}$:
${\widetilde{d}}_{n}(\nu_{0},\kappa)=\sum_{s=0}^{n}{\widetilde{k}}_{n-s}(\nu_{0}+s)d_{s}(\nu_{0};\kappa).$ (3)
Analogously, when we use the relation (2a) in the expansion (1b), and take into account the notation (1a), we obtain the relations between the coefficients $d_{n}$ and ${\widetilde{d}}_{k}$
$d_{n}(\nu_{0},\kappa)=\sum_{s=0}^{n}k_{n-s}(\nu_{0}+s){\widetilde{d}}_{s}(\nu_{0};\kappa).$ (4)
The (generalised) logarithmic derivatives obey the differential (recursive) relation GCAK
$\frac{d}{d\ln\kappa}{\widetilde{a}}_{\nu}(\kappa Q^{2})=(-\beta_{0})\nu{\widetilde{a}}_{\nu+1}(\kappa Q^{2}),$ (5)
where $\beta_{0}=(11-2n_{f}/3)/4$ is the one-loop beta coefficient of the renormalisation group equation (RGE) of the running $a(\kappa Q^{2})$
$\frac{d{a}(\mu^{2})}{d\ln\mu^{2}}$ $=$ $-\beta_{0}a(\mu^{2})^{2}-\beta_{1}a(\mu^{2})^{3}-\beta_{2}a(\mu^{2})^{4}-\beta_{3}a(\mu^{2})^{5}$-$\ldots$ (6a)
$=$ $-\beta_{0}a(\mu^{2})^{2}\left[1+c_{1}a(\mu^{2})+c_{2}a(\mu^{2})^{2}+c_{3}a(\mu^{2})^{3}+\ldots\right].$ (6b)
In the mass-independent schemes (such as $\overline{\rm MS}$ or MiniMOM), the first two coefficients ($\beta_{0}$, $\beta_{1}$) are universal, i.e., scheme-independent. [^4] We have: $\beta_{0}=(11-2n_{f}/3)/4$ and $\beta_{1}=(102-38n_{f}/3)/16$, where $n_{f}$ is the number of active massless quark flavours. At low $|Q^{2}|\lesssim 1\ {\rm GeV}^{2}$ we have $n_{f}=3$. The coefficients $\beta_{j}$ (or $c_{j}\equiv\beta_{j}/\beta_{0}$) with $j=2,3,\ldots$ then define the renormalisation scheme. Since ${\cal D}(Q^{2})$ is an observable, it is independent of the renormalisation scale parameter $\kappa$ ($\equiv\mu^{2}/Q^{2}$), i.e., $(d/d\ln\kappa){\cal D}(Q^{2})=0$. Applying this condition to the expansion (1b) for ${\cal D}(Q^{2})$ and using the recursive relation (5) leads to
$\frac{d}{d\ln\kappa}{\cal D}(Q^{2})$ $=$ ${\widetilde{a}}_{\nu_{0}}(\kappa Q^{2})\left[\frac{d}{d\ln\kappa}{\widetilde{d}}_{0}(\nu_{0},\kappa)\right]+\sum_{n=1}^{\infty}{\widetilde{a}}_{\nu_{0}+n}(\kappa Q^{2})\left[(-\beta_{0})(\nu_{0}+n-1){\widetilde{d}}_{n-1}(\nu_{0};\kappa)+\frac{d}{d\ln\kappa}{\widetilde{d}}_{n}(\nu_{0};\kappa)\right],$ (7)
which implies that the coefficient at each ${\widetilde{a}}_{\nu_{0}+n}$ in this expression must be zero
$\frac{d}{d\ln\kappa}{\widetilde{d}}_{n}(\nu_{0},\kappa)=\beta_{0}(n+\nu_{0}-1){\widetilde{d}}_{n-1}(\nu_{0};\kappa)\qquad(n=1,2,\ldots),$ (8)
and for $n=0$ we get $(d/d\ln\kappa){\widetilde{d}}_{0}(\nu_{0},\kappa)=0$, which allows us to choose ${\widetilde{d}}_{0}(\nu_{0};\kappa)=1$, i.e., the canonical choice. In Eq. (8) we see that the $\kappa$-derivative of the coefficients ${\widetilde{d}}_{n}(\nu_{0};\kappa)$ yields a factor on the right-hand side (RHS) where the index $\nu_{0}$ explicitly appears.
The next step is to improve the evaluation of the observable ${\cal D}(Q^{2})$, from using the simple (truncated) expansions Eqs. (1) to using a renormalon-motivated resummation procedure. This is done by generalising the resummation method which has been developed for the integer case of $\nu_{0}$ ($\nu_{0}=1$) in Ref. renmod to the case of noninteger $\nu_{0}$. This suggests that we now consider a modified quantity, with modified coefficients ${\widetilde{d}}_{n}(\nu_{0};\kappa)\mapsto{\widetilde{d}}_{n}(1;\kappa)$ where the resulting factor on the RHS of Eq. (8) becomes $(n+\nu_{0}-1)\mapsto n$. This is achieved with the following rescaling of the original coefficients ${\widetilde{d}}_{n}(\nu_{0};\kappa)$ that appear in the expansion of (1b):
${\widetilde{d}}_{n}(1;\kappa)\equiv\frac{\Gamma(\nu_{0})\Gamma(1+n)}{\Gamma(\nu_{0}+n)}{\widetilde{d}}_{n}(\nu_{0};\kappa).$ (9)
When we use this relation in the recursive relations (8), we obtain indeed
$\frac{d}{d\ln\kappa}{\widetilde{d}}_{n}(1,\kappa)=\beta_{0}n{\widetilde{d}}_{n-1}(1;\kappa)\qquad(n=1,2,\ldots),$ (10)
and ${\widetilde{d}}_{0}(1;\kappa)=1$.
If we now define a new auxiliary quantity ${{\cal D}^{(1)}}(Q^{2})$ whose perturbation expansion in logarithmic derivatives involves these rescaled coefficients
${{\cal D}^{(1)}}(Q^{2})\equiv\sum_{n=0}^{\infty}{\widetilde{d}}_{n}(1;\kappa){\widetilde{a}}_{n+1}(\kappa Q^{2}),$ (11)
we can see immediately that this quantity is a quasiobservable, i.e., it is independent of the renormalisation scale parameter $\kappa$: $(d/d\ln\kappa){{\cal D}^{(1)}}(Q^{2})=0$. We note that the expansion (11) starts with ${\widetilde{a}}_{1}(\kappa Q^{2})$ which is identical with $a(\kappa Q^{2})$. The logarithmic derivatives ${\widetilde{a}}_{n+1}(\kappa Q^{2})$ appearing in the expansion (11) have now integer indices $n+1$, and are literally logarithmic derivatives of $a(\kappa Q^{2})$ (cf. footnote 1)
${\widetilde{a}}_{n+1}(Q^{2})\equiv\frac{(-1)^{n}}{n!\beta_{0}^{n}}\left(\frac{d}{d\ln Q^{2}}\right)^{n}a(Q^{2}).$ (12)
Following Ref. renmod , we now define the modified Borel transform $\widetilde{B}$} of this quantity ${{\cal D}^{(1)}}(Q^{2})$
${\widetilde{B}}[{{\cal D}^{(1)}}](u;\kappa)\equiv\sum_{n=0}^{\infty}\frac{{\widetilde{d}}_{n}(1;\kappa)}{n!\beta_{0}^{n}}u^{n},$ (13)
for which we know, from our treatment of the $\nu_{0}=1$ case renmod (cf. also BSRpQCD ), that it has the following simple (one-loop type) $\kappa$-dependence: [^5] This property is exact and is the direct consequence of the definition (13) and the recursion relations (10), where the latter are the consequence of the $\kappa$-independence of ${{\cal D}^{(1)}}(Q^{2})$ $\Leftrightarrow$ $\kappa$-independence of the original quantity ${\cal D}(Q^{2})$ Eqs. (1a)-(1b)].
${\widetilde{B}}[{{\cal D}^{(1)}}](u;\kappa)=\kappa^{u}{\widetilde{B}}[{{\cal D}^{(1)}}](u),$ (14)
and that we can resum the quantity ${{\cal D}^{(1)}}(Q^{2})$ with a characteristic function $F_{{{\cal D}^{(1)}}}(t)$
${{\cal D}^{(1)}}(Q^{2})_{\rm res.}=\int_{0}^{\infty}\frac{dt}{t}F_{{{\cal D}^{(1)}}}(t)a(tQ^{2}),$ (15)
where the characteristic function is the inverse Mellin transformation of the modified Borel ${\widetilde{B}}[{{\cal D}^{(1)}}]$
$F_{{{\cal D}^{(1)}}}(t)=\frac{1}{2\pi i}\int_{u_{0}-i\infty}^{u_{0}+\infty}du{\widetilde{B}}[{{\cal D}^{(1)}}](u)t^{u}.$ (16)
Here, $u_{0}$ is zero, or any real number closer to zero than the first renormalon singularity of ${\widetilde{B}}[{{\cal D}^{(1)}}](u)$.
We can check that the resummation (15) is correct, if we Taylor-expand $a(tQ^{2})$ there around the point $\kappa Q^{2}$ [in fact: around the point $\ln(\kappa Q^{2})$]
$a(tQ^{2})=\sum_{n=0}^{\infty}(-\beta_{0})^{n}\ln^{n}\left(\frac{t}{\kappa}\right){\widetilde{a}}_{n+1}(\kappa Q^{2}),$ (17)
insert this expansion in Eq. (15), exchange the order of the summation and integration, and require that we obtain the expansion (11) of ${\cal D}^{(1)}$}, which then gives at each ${\widetilde{a}}_{n+1}(\kappa Q^{2})$ the condition
${\widetilde{d}}_{n}(1;\kappa)=(-\beta_{0})^{n}\int_{0}^{\infty}\frac{dt}{t}F_{{{\cal D}^{(1)}}}(t)\ln^{n}\left(\frac{t}{\kappa}\right)\qquad(n=0,1,2,\ldots).$ (18)
We then multiply each of these sum rules by $u^{n}/(n!\beta_{0}^{n})$ and sum over $n$; we must obtain in this way the modified Borel ${\widetilde{B}}[{{\cal D}^{(1)}}](u;\kappa)$ Eq. (13); this then implies [^6] We note: $\sum_{n=0}^{\infty}(-1)^{n}w^{n}/n!=e^{-w}$, where we have in our case $w=u\ln(t/\kappa)$.
${\widetilde{B}}[{{\cal D}^{(1)}}](u;\kappa)=\kappa^{u}\int_{0}^{\infty}\frac{dt}{t}F_{{{\cal D}^{(1)}}}(t)t^{-u},$ (19)
where we have expressed the obtained exponential as $\exp(-u\ln(t/\kappa))=\kappa^{u}t^{-u}$. The $\kappa$-dependence in this expression is consistent with the property (14) that reflects the $\kappa$-independence of ${{\cal D}^{(1)}}(Q^{2})$ and ${\cal D}(Q^{2})$. When $\kappa=1$, we thus obtain
${\widetilde{B}}[{{\cal D}^{(1)}}](u)=\int_{0}^{\infty}\frac{dt}{t}F_{{{\cal D}^{(1)}}}(t)t^{-u},$ (20)
which means that ${\widetilde{B}}[{{\cal D}^{(1)}}](u)$ is the Mellin transform of the characteristic function $F_{{{\cal D}^{(1)}}}(t)$ that appeared in the resummation of ${{\cal D}^{(1)}}(Q^{2})$, Eq. (15). Thus the (sought) characteristic function for the resummation of ${{\cal D}^{(1)}}(Q^{2})$ is the inverse Mellin of the modified Borel ${\widetilde{B}}[{{\cal D}^{(1)}}](u)$, i.e., Eq. (15) is proven. [^7] A similar formalism of resummation, applicable to observables with expansion in powers of the perturbative running coupling in the one-loop approximation, $a_{\rm(1\ell)}(Q^{2})$, was developed first by Neubert Neubert and later applied also in NC . The formalism of renmod , however, works when the running coupling $a(Q^{2})$ is taken at an arbitrary loop-level. We note that, at the one-loop level, the powers and the logarithmic derivatives (for integer $n$) coincide, $a_{\rm(1\ell)}^{n}={\widetilde{a}}_{n}^{\rm(1\ell)}$.
Until now, we presented and explained relations for ${{\cal D}^{(1)}}(Q^{2})$, i.e., for the case $\nu_{0}\mapsto 1$, which is a recapitulation of the results already obtained in renmod and BSRpQCD ; BSRAQCD . Now we proceed to the resummation of the original observable ${\cal D}(Q^{2})$ which has $\nu_{0}\not=1$ (i.e., $\nu_{0}$ in general noninteger), i.e., whose expansion is given in Eqs. (1). It turns out that this step consists in simply replacing, in the resummation Eq. (15) for the auxiliary quasiobservable ${{\cal D}^{(1)}}(Q^{2})$, the factor $a(tQ^{2})$ by ${\widetilde{a}}_{\nu_{0}}(tQ^{2})$:
Theorem 1:
The resummation of ${\cal D}(Q^{2})$ observable as characterised by its expansions Eqs. (1a)-(1b) yields
${\cal D}(Q^{2})_{\rm res.}=\int_{0}^{\infty}\frac{dt}{t}F_{{{\cal D}^{(1)}}}(t){\widetilde{a}}_{\nu_{0}}(tQ^{2}),$ (21)
where $F_{{{\cal D}^{(1)}}}(t)$ is the inverse Mellin, Eq. (16), of the auxiliary quasiobservable ${{\cal D}^{(1)}}(Q^{2})$ which is defined by the perturbation expansion (11) with the rescaled coefficients $d_{n}(1;\kappa)$ defined via relations (9). The generalised logarithmic derivative ${\widetilde{a}}_{\nu_{0}}(tQ^{2})$ in the result Eq. (21) is a linear combination of powers $a(tQ^{2})^{\nu_{0}+n}$ appearing in Eq. (2a) where the general coefficients $k_{m}(\nu_{0})$ were obtained in GCAK .
Proof:
In order to prove the resummation formula (21), we perform the Taylor-expansion of ${\widetilde{a}}_{\nu_{0}}(tQ^{2})$ around $\kappa Q^{2}$ [analogous to the expansion Eq. (17)], by using the recursive relations (5)
${\widetilde{a}}_{\nu_{0}}(tQ^{2})=\sum_{n=0}^{\infty}(-\beta_{0})^{n}\ln^{n}\left(\frac{t}{\kappa}\right)\frac{\Gamma(\nu_{0}+n)}{\Gamma(\nu_{0})\Gamma(n+1)}{\widetilde{a}}_{\nu_{0}+n}(\kappa Q^{2}),$ (22)
We use this expansion on the RHS of Eq. (21), and exchange the order of the summation and integration. The term at ${\widetilde{a}}_{\nu_{0}+n}(\kappa Q^{2})$ must then give the coefficient ${\widetilde{d}}_{n}(\nu_{0};\kappa)$ of the expansion (1b)
${\widetilde{d}}_{n}(\nu_{0};\kappa)$ $=$ $\frac{\Gamma(\nu_{0}+n)}{\Gamma(\nu_{0})\Gamma(n+1)}(-\beta_{0})^{n}\int_{0}^{\infty}\frac{dt}{t}F_{{{\cal D}^{(1)}}}(t)\ln^{n}\left(\frac{t}{\kappa}\right)\Rightarrow$ (23a)
${\widetilde{d}}_{n}(1;\kappa)\left[\equiv\frac{\Gamma(\nu_{0})\Gamma(n+1)}{\Gamma(\nu_{0}+n)}{\widetilde{d}}_{n}(\nu_{0};\kappa)\right]$ $=$ $(-\beta_{0})^{n}\int_{0}^{\infty}\frac{dt}{t}F_{{{\cal D}^{(1)}}}(t)\ln^{n}\left(\frac{t}{\kappa}\right)\qquad(n=0,1,2,\ldots)$ (23b)
The form (23b) is indeed true, by the construction of the characteristic function $F_{{{\cal D}^{(1)}}}(t)$, cf. the sum rules (18). This then proves that the result (21) is the correct resummation because it implies the original expansion (1b) of the considered observable ${\cal D}(Q^{2})$.
In practice, in order to obtain the characteristic function $F_{{{\cal D}^{(1)}}}(t)$ needed in the resummation Eq. (21) of the spacelike QCD observable ${\cal D}(Q^{2})$ [whose expansions are in Eqs. (1)], we will need to know the modified Borel transform ${\widetilde{B}}[{{\cal D}^{(1)}}](u)$ of the auxiliary quantity ${{\cal D}^{(1)}}(Q^{2})$. Therefore, we face the following problem: if we know the renormalon structure of this observable ${\cal D}(Q^{2})$, [^8] For a review of renormalon physics, see Renormalons . i.e., the behaviour of the (usual) Borel transform $B^{(\nu_{0})}[{\cal D}](u)$ of $\cal D$}, what is the corresponding behaviour of the modified Borel ${\widetilde{B}}^{(\nu_{0})}[{\cal D}](u)$ of $\cal D$}, and of the modified Borel ${\widetilde{B}}[{{\cal D}^{(1)}}](u)$? Namely, we need the behaviour of the latter quantity in order to get the characteristic function $F_{{{\cal D}^{(1)}}}(t)$ that is the inverse Mellin of ${\widetilde{B}}[{{\cal D}^{(1)}}](u)$ Eq. (16), i.e., the characteristic function appearing also in the resummation (21) of the obervable ${\cal D}(Q^{2})$. We will adopt for the observable ${\cal D}(Q^{2})$, Eqs. (1), the following definitions of the Borel and modified Borel transforms:
$B^{(\nu_{0})}[{\cal D}](u;\kappa)$ $\equiv$ $\sum_{n=0}^{\infty}\frac{d_{n}(\nu_{0};\kappa)}{n!\beta_{0}^{n}}u^{n},$ (24a)
${\widetilde{B}}^{(\nu_{0})}[{\cal D}](u;\kappa)$ $\equiv$ $\sum_{n=0}^{\infty}\frac{{\widetilde{d}}_{n}(\nu_{0};\kappa)}{n!\beta_{0}^{n}}u^{n},$ (24b)
and the definition of ${\widetilde{B}}[{{\cal D}^{(1)}}](u;\kappa)$ is given in Eq. (13). [^9] The modified Borel transforms of the quantities with $\nu_{0}=1$ (such as ${\cal D}^{(1)}$}) will be denoted, for simplicity, without the superscript ’$(1)$’ at $\widetilde{B}$}.
The above problem is addressed by the following Theorems 2 and 3. We first assume that the renormalon structures are dominated by an infrared (IR) renormalon at $u=p$ ($p>0$), and then comment on the modification for the case of an ultraviolet (UV) renormalon at u=pu=-p.
Theorem 2:
If the (IR) renormalon structure of the Borel transform, Eq. (24a), of the observable ${\cal D}(Q^{2})$ [Eq. (1)] has the following form:
$B^{(\nu_{0})}[{\cal D}](u;\kappa)=\frac{{\cal K}(\kappa)}{(p-u)^{s_{0}}}\left[1+{\cal O}((p-u))\right],$ (25)
where $p>0$ (in practice: $p=1/2,1,3/2,\ldots$), then the modified Borel transform (24b) of this observable is
${\widetilde{B}}^{(\nu_{0})}[{\cal D}](u;\kappa)=\frac{{\widetilde{\cal K}}(\kappa)}{(p-u)^{{\widetilde{s}}_{0}}}\left[1+{\cal O}((p-u))\right],$ (26)
where the two indices are related by
$s_{0}={\widetilde{s}}_{0}+p\frac{\beta_{1}}{\beta_{0}^{2}},$ (27)
where $\beta_{0}$ and $\beta_{1}$ are the two leading coefficients in the $\beta$-function (6).
The proof of this theorem is relatively long, and we refer to Appendix A for the proof. The residues ${\cal K}(\kappa)$ and ${\widetilde{\cal K}}(\kappa)$ are also related, but we will not need their relation in our application. [^10] An approximate relation between these two residues can be inferred by comparing Eq. (105b) with Eqs. (120) and (129).
If the renormalon is at u=pu=-p (i.e., UV renormalon), then we replace in Eqs. (25)-(26): $(p-u)\mapsto(p+u)$; and in the relation (27): $p\mapsto-p$. The proof can be repeated in this case (with: $p\mapsto-p$)
The validity of this Theorem was checked numerically for quantities with $\nu_{0}=1$ and several integer ${\widetilde{s}}_{0}$ and integer $p$ in Ref. renmod ; and for noninteger ${\widetilde{s}}_{0}=0.778$ and ${\widetilde{s}}_{0}=0.375$ (with $p=3$ and $\nu_{0}=1$) in Ref. JPGour . [^11] At the time, the general validity of Theorem 2 (i.e., for noninteger $\nu_{0}$, ${\widetilde{s}}_{0}$ and $p$) was not known or used in those references.
Theorem 3:
If the modified Borel transform ${\widetilde{B}}^{(\nu_{0})}[{\cal D}](u;\kappa)$ Eq. (24b) of the observable ${\cal D}(Q^{2})$ Eq. (1) has the renormalon form as given in Eq. (26), then the modified Borel transform ${\widetilde{B}}[{{\cal D}^{(1)}}](u;\kappa)$ Eq. (13) of the auxiliary quasiobservable ${{\cal D}^{(1)}}(Q^{2})$ Eq. (11) [cf. Eq. (9)] is
${\widetilde{B}}[{{\cal D}^{(1)}}](u;\kappa)=\frac{{{\widetilde{\cal K}}^{(1)}}(\kappa)}{(p-u)^{{\widetilde{s}}_{0}-\nu_{0}+1}}\left[1+{\cal O}((p-u))\right],$ (28)
Proof:
We will use the general expansion ($s$ can be noninteger)
$(p-u)^{-s}=p^{-s}\sum_{n=0}^{\infty}\frac{\Gamma(s+n)}{\Gamma(s)n!}\left(\frac{u}{p}\right)^{n},$ (29)
and the asymptotic formula for the $\Gamma$ function (for large $n$)
$\frac{\Gamma(s+n)}{n!}=n^{s-1}\left[1+{\cal O}(1/n)\right].$ (30)
Using this, the asymptotic expansion of Eq. (26) is
${\widetilde{B}}^{(\nu_{0})}[{\cal D}](u;\kappa)={\widetilde{\cal K}}(\kappa)\frac{p^{-{\widetilde{s}}_{0}}}{\Gamma({\widetilde{s}}_{0})}\sum_{n}\frac{n^{{\widetilde{s}}_{0}-1}}{p^{n}}\left(1+{\cal O}(1/n)\right)u^{n},$ (31)
which implies the following asymptotic behavior of the corresponding coefficients ${\widetilde{d}}_{n}(\nu_{0};\kappa)$:
${\widetilde{d}}_{n}(\nu_{0};\kappa)$ $=$ ${\widetilde{\cal K}}(\kappa)\frac{p^{-{\widetilde{s}}_{0}}}{\Gamma({\widetilde{s}}_{0})}n!\;n^{{\widetilde{s}}_{0}-1}\left(\frac{\beta_{0}}{p}\right)^{n}\left(1+{\cal O}(1/n)\right).$ (32)
The asymptotic behavior of the corresponding modified coefficients ${\widetilde{d}}_{n}(1;\kappa)$ [as defined by Eq. (9)] is then
${\widetilde{d}}_{n}(1;\kappa)$ $=$ $\frac{\Gamma(\nu_{0})n!}{\Gamma(\nu_{0}+n)}{\widetilde{\cal K}}(\kappa)\frac{p^{-{\widetilde{s}}_{0}}}{\Gamma({\widetilde{s}}_{0})}n!\;n^{{\widetilde{s}}_{0}-1}\left(\frac{\beta_{0}}{p}\right)^{n}\left(1+{\cal O}(1/n)\right)$ (33a)
$=$ ${\widetilde{\cal K}}(\kappa)\frac{\Gamma(\nu_{0})p^{-{\widetilde{s}}_{0}}}{\Gamma({\widetilde{s}}_{0})}n!\;n^{({\widetilde{s}}_{0}-\nu_{0}+1)-1}\left(\frac{\beta_{0}}{p}\right)^{n}\left(1+{\cal O}(1/n)\right),$ (33b)
where we used in the last identity (33b) the asymptotic formula (30). When we compare the expression (33b) with (32), we see that the coefficients ${\widetilde{d}}_{n}(1;\kappa)$ give in the Borel transform the same structure as ${\widetilde{d}}_{n}(\nu_{0};\kappa)$, except that ${\widetilde{s}}_{0}\mapsto{\widetilde{s}}_{0}-\nu_{0}+1$ in the exponent of $n$. This then implies, that the renormalon structure in ${\widetilde{B}}[{{\cal D}^{(1)}}](u;\kappa)$ is the same as in ${\widetilde{B}}^{(\nu_{0})}[{\cal D}](u;\kappa)$, when we replace in the index ${\widetilde{s}}_{0}\mapsto{\widetilde{s}}_{0}-\nu_{0}+1$. This then proves Theorem 3.121212 The relative corrections ${\cal O}(1/n)$ in Eq. (33b) correspond to the relative corrections ${\cal O}(p-u)$ in the expression (28), because $\Gamma((s-1)+n)=(\Gamma(s+n)/n)$ $\times(1+{\cal O}(1/n))$.
If the renormalon is at u=pu=-p (UV renormalon), then we replace in the theorem Eq. (28): $(p-u)\mapsto(p+u)$, and the proof can be repeated (with: $p\mapsto-p$).
In order to perform in practice the resummation Eq. (21) for general spacelike QCD observables ${\cal D}(Q^{2})$, Eqs. (1), we need to obtain the associated characteristic function $F_{{{\cal D}^{(1)}}}(t)$, Eq. (16), that is the inverse Mellin transform of the modified Borel ${\widetilde{B}}[{{\cal D}^{(1)}}](u)$. The following theorem enables us to evaluate the characteristic function $F_{{{\cal D}^{(1)}}}(t)$ if we know the renormalon structure of the Borel $B^{(\nu_{0})}[{\cal D}](u)$ Eq. (25) [and thus ${\widetilde{B}}[{{\cal D}^{(1)}}](u)$ by Theorems 2 and 3]:
Theorem 4:
If the modified Borel transform ${\widetilde{B}}[{{\cal D}^{(1)}}](u;\kappa)$ (with $\kappa=1$) has the (IR) renormalon contribution
${\widetilde{B}}[{{\cal D}^{(1)}}](u)_{(p,{\widetilde{s}})}=\frac{\pi}{(p-u)^{{\widetilde{s}}}}$, (34)
where $p>0$ and $0<{\widetilde{s}}\leq 1$, then the corresponding inverse Mellin transform Eq. (16), i.e., the corresponding characteristic function $F_{{{\cal D}^{(1)}}}(t)_{(p,{\widetilde{s}})}$, is
$F_{{{\cal D}^{(1)}}}(t)_{(p,{\widetilde{s}})}=\Theta(1-t)\pi\frac{t^{p}}{\Gamma({\widetilde{s}})(-\ln t)^{1-{\widetilde{s}}}}$, (35)
where $\Theta(1-t)$ is the Heaviside function (i.e., it is unity for $0\leq t\leq 1$, and is zero for $t>1$).
We refer for a formal proof of this theorem to Appendix B. This theorem, together with Theorems 2 and 3, implies that the characteristic function $F_{{{\cal D}^{(1)}}}(t)$ appearing in the resummation (21) of the spacelike observable ${\cal D}(Q^{2})$ Eqs. (1) with the IR renormalon structure (25) is
$F_{{{\cal D}^{(1)}}}(t)=\Theta(1-t){{\widetilde{\cal K}}^{(1)}}\frac{t^{p}}{\Gamma({\widetilde{s}})(-\ln t)^{1-{\widetilde{s}}}}\qquad{\rm with:}\;\;{\widetilde{s}}=s_{0}-p\frac{\beta_{1}}{\beta_{0}^{2}}-\nu_{0}+1,$ (36)
where ${\widetilde{\cal K}}^{(1)}$} is a constant.131313 We took here $\kappa=1$. We note that then: ${{\widetilde{\cal K}}^{(1)}}(\kappa)=\kappa^{p}{{\widetilde{\cal K}}^{(1)}}$, because $\kappa^{u}=\kappa^{p}[1+{\cal O}((p-u))]$. We point out that in our approach we take into account only the leading (i.e., the most singular) renormalon contributions, i.e., we neglect the relative corrections ${\cal O}((p-u)^{1})$ in the (modified) Borel transforms. According to Eq. (36), these relative corrections then contribute to the characteristic function $F_{{{\cal D}^{(1)}}}(t)$ the relative corrections ${\cal O}(1/|\ln t|)$ (we note that $1/|\ln t|<1$ for $0\leq t<1/e$ ($\approx 0.37$).
When, however, we have in the modified Borel transform ${\widetilde{B}}[{{\cal D}^{(1)}}](u)$ the effects beyond the leading term (34) contained in the rescaling factor $\exp({\widetilde{K}}_{e}u)$ [reflecting possible redefinition of the momentum scale, according to Eq. (14)]
${\widetilde{B}}[{{\cal D}^{(1)}}](u)_{(p,{\widetilde{s}},{\widetilde{K}}_{e})}=\frac{\pi\exp({\widetilde{K}}_{e}u)}{(p-u)^{{\widetilde{s}}}}$, (37)
then it is straightforward to check that the inverse Mellin (16) implies for the characteristic function in such a case
$F_{{{\cal D}^{(1)}}}(t)_{(p,{\widetilde{s}},{\widetilde{K}}_{e})}=F_{{{\cal D}^{(1)}}}(te^{{\widetilde{K}}_{e}})_{(p,{\widetilde{s}})}$, (38)
where $F_{{{\cal D}^{(1)}}}(t)_{(p,{\widetilde{s}})}$ is given in Eq. (35). This then implies, after a simple redefinition of the integration variable $t$, that the resummation expressions (15), (21), in the case of the modified Borel (37), have the form
${{\cal D}^{(1)}}(Q^{2})_{\rm res.}$ $=$ $\int_{0}^{\infty}\frac{dt}{t}F_{{{\cal D}^{(1)}}}(t)_{(p,{\widetilde{s}})}a(te^{-{\widetilde{K}}_{e}}Q^{2}),$ (39a)
${\cal D}(Q^{2})_{\rm res.}$ $=$ $\int_{0}^{\infty}\frac{dt}{t}F_{{{\cal D}^{(1)}}}(t)_{(p,{\widetilde{s}})}{\widetilde{a}}_{\nu_{0}}(te^{-{\widetilde{K}}_{e}}Q^{2}).$ (39b)
III Resummation for timelike observable ${\cal F}(\sigma)$
We will now consider the timelike observable ${{\cal F}}(\sigma)$ ($\sigma>0$) that is associated with the spacelike observable ${\cal D}(Q^{2})$ of Eqs. (1), and the auxiliary timelike quasiobservable ${{{\cal F}^{(1)}}}(\sigma)$ that is associated with the auxiliary spacelike quasiobservable ${{\cal D}^{(1)}}(Q^{2})$ of Eq. (11) [in conjunction with Eq. (9)]
${\cal F}(\sigma)$ $=$ $\frac{1}{2\pi}\int_{-\pi}^{+\pi}d\phi\;{\cal D}(\sigma e^{i\phi}),$ (40a)
${{\cal F}^{(1)}}(\sigma)$ $=$ $\frac{1}{2\pi}\int_{-\pi}^{+\pi}d\phi\;{{\cal D}^{(1)}}(\sigma e^{i\phi}).$ (40b)
We recall that the inverse relations have the somewhat more familiar form
${\cal D}(Q^{2})$ $=$ $Q^{2}\int_{0}^{\infty}\frac{d\sigma{\cal F}(\sigma)}{(\sigma+Q^{2})^{2}},$ (41a)
${{\cal D}^{(1)}}(Q^{2})$ $=$ $Q^{2}\int_{0}^{\infty}\frac{d\sigma{{\cal F}^{(1)}}(\sigma)}{(\sigma+Q^{2})^{2}}$ (41b)
The corresponding expansions in generalised logarithmic derivatives for these timelike quantities, analogous to the expansions (1b) and (11), are denoted analogously as
${\cal F}(\sigma)$ $=$ $\sum_{n=0}^{\infty}{\widetilde{f}}_{n}(\nu_{0};\kappa){\widetilde{a}}_{\nu_{0}+n}(\kappa\sigma),$ (42a)
${{\cal F}^{(1)}}(\sigma)$ $=$ $\sum_{n=0}^{\infty}{\widetilde{f}}_{n}(1;\kappa){\widetilde{a}}_{1+n}(\kappa\sigma).$ (42b)
First we will prove the following theorem which relates the modified Borel transform $\widetilde{B}$} of the quantity ${{\cal F}^{(1)}}(\sigma)$ with $\widetilde{B}$} of ${{\cal D}^{(1)}}(Q^{2})$:
Theorem 5: We have
${\widetilde{B}}[{{\cal D}^{(1)}}](u;\kappa)\frac{\sin(\pi u)}{\pi u}={\widetilde{B}}[{{\cal F}^{(1)}}](u;\kappa),$ (43)
where ${\widetilde{B}}[{{\cal D}^{(1)}}](u;\kappa)$ was defined through its expansion in Eq. (13), and ${\widetilde{B}}[{{\cal F}^{(1)}}](u\kappa)$ is defined via the corresponding expansion
${\widetilde{B}}[{{\cal F}^{(1)}}](u;\kappa)=\sum_{n=0}^{\infty}\frac{{\widetilde{f}}_{n}(1;\kappa)}{n!\beta_{0}^{n}}u^{n}.$ (44)
Proof:
In this proof we follow closely the steps applied in Appendix A of our previous paper EPJC81 [for $\delta_{x^{n}}^{(d)}$ there, with $n=0$].
The idea is to operationally replace the logarithmic-derivative couplings ${\widetilde{a}}_{1+n}(\kappa Q^{2})$ and ${\widetilde{a}}_{1+n}(\kappa\sigma)$ in the expansions (11) and (42b) by the simple powers $a_{\rm(1\ell)}(\kappa Q^{2})^{1+n}$ and $a_{\rm(1\ell)}(\kappa\sigma)^{1+n}$, where $a_{\rm(1\ell)}(Q^{2})$ is the one-loop running coupling. For the purpose of the present proof, this is legitimate, because the momentum dependence of those logarithmic-derivative couplings is exactly the same as that of one-loop coupling powers. This can be seen directly from the relation (5) (when $\nu=1+n$). [^14] We point out, however, that in our approach the logarithmic derivatives are considered at any loop level., [^15] As a consequence, the scale-dependence relation (22) remains valid also when we replace there everywhere ${\widetilde{a}}_{\nu_{0}+n}\mapsto a_{\rm(1\ell)}^{\nu_{0}+n}$, and $\nu_{0}$ can be either integer or noninteger. We can also interpret this replacement by the fact that in the one-loop approximation ($\beta_{j}=0$ for $j\geq 1$) we have clearly ${\widetilde{a}}_{1+n}=a^{n+1}$.
Following this approach, we introduce expansions in powers of the one-loop coupling analogous to the exact expansions (11) and (42b)
${{\cal D}^{(1)}}_{\rm(1\ell,pow.)}(Q^{2})$ $=$ $\sum_{n=0}^{\infty}{\widetilde{d}}_{n}(1;\kappa)a_{\rm(1\ell)}(\kappa Q^{2})^{1+n},$ (45a)
${{\cal F}^{(1)}}_{\rm(1\ell,pow.)}(\sigma)$ $=$ $\sum_{n=0}^{\infty}{\widetilde{f}}_{n}(1;\kappa)a_{\rm(1\ell)}(\kappa\sigma)^{1+n}.$ (45b)
We point out that these power expansions contain exactly the same coefficients ${\widetilde{d}}_{n}(1;\kappa)$ and ${\widetilde{f}}_{n}(1;\kappa)$ of the (exact) expansions (11) and (42b) of the quantities ${{\cal D}^{(1)}}(Q^{2})$ and ${{\cal F}^{(1)}}(\sigma)$. The quantities (45a) and (45b) are exactly independent of the renormalisation scale parameter $\kappa$ [as are ${{\cal D}^{(1)}}(Q^{2})$ and ${{\cal F}^{(1)}}(\sigma)$]. Now, the usual formal Borel transforms of $B[{{\cal D}^{(1)}}_{\rm(1\ell,pow.)}]$ and $B[{{\cal F}^{(1)}}_{\rm(1\ell,pow.)}]$ are
$B[{{\cal D}^{(1)}}_{\rm(1\ell,pow.)}](u;\kappa)$ $=$ $\sum_{n=0}^{\infty}\frac{{\widetilde{d}}_{n}(1;\kappa)}{n!\beta_{0}^{n}}u^{n}\quad\left[={\widetilde{B}}[{{\cal D}^{(1)}}](u;\kappa)\right],$ (46a)
$B[{{\cal F}^{(1)}}_{\rm(1\ell,pow.)}](u;\kappa)$ $=$ $\sum_{n=0}^{\infty}\frac{{\widetilde{f}}_{n}(1;\kappa)}{n!\beta_{0}^{n}}u^{n}\quad\left[={\widetilde{B}}[{{\cal F}^{(1)}}](u;\kappa)\right],$ (46b)
which, as indicated, are identical to the modified Borel transforms ${\widetilde{B}}[{{\cal D}^{(1)}}](u;\kappa)$ Eq. (13) and ${\widetilde{B}}[{{\cal F}^{(1)}}](u;\kappa)$ Eq. (44), respectively. The inverse Borel transformation of the usual Borel $B[{{\cal D}^{(1)}}_{\rm(1\ell,pow.)}](u;\kappa)$ is then
${{\cal D}^{(1)}}_{\rm(1\ell,pow.)}(Q^{2})=\frac{1}{\beta_{0}}\int_{0}^{\infty}du\;\exp\left(-\frac{u}{\beta_{0}a_{\rm(1\ell)}(\kappa Q^{2})}\right){\widetilde{B}}[{{\cal D}^{(1)}}](u;\kappa).$ (47)
This then implies that the corresponding timelike quantity ${{\cal F}^{(1)}}_{\rm(1\ell,pow.)}(\sigma)$ is [cf. Eqs. (40b) and (45b)]
${{\cal F}^{(1)}}_{\rm(1\ell,pow.)}(\sigma)$ $=$ $\frac{1}{2\pi}\int_{-\pi}^{+\pi}d\phi\;{{\cal D}^{(1)}}_{\rm(1\ell,pow.)}(\sigma e^{i\phi})$ (48a)
$=$ $\frac{1}{2\pi\beta_{0}}\int_{-\pi}^{+\pi}d\phi\;\int_{0}^{\infty}du\;\exp\left(-\frac{u}{\beta_{0}a_{\rm(1\ell)}(\kappa\sigma e^{i\phi})}\right){\widetilde{B}}[{{\cal D}^{(1)}}](u;\kappa)$ (48b)
$=$ $\frac{1}{2\pi\beta_{0}}\int_{0}^{\infty}\;du\exp\left(-\frac{u}{\beta_{0}a_{\rm(1\ell)}(\kappa\sigma)}\right){\widetilde{B}}[{{\cal D}^{(1)}}](u;\kappa)\;\int_{-\pi}^{+\pi}d\phi\exp\left(-\frac{i\beta_{0}u}{\beta_{0}}\phi\right),$ (48c)
where in the last identity we exchanged the order of integrations and used the one-loop RGE-running relation
$\frac{1}{a_{\rm(1\ell)}(\kappa\sigma e^{i\phi})}=\frac{1}{a_{\rm(1\ell)}(\kappa\sigma)}+i\beta_{0}\phi.$ (49)
The integral over $\phi$ in Eq. (48c) is trivial, equal to $2\sin(\pi u)/u$, leading to
${{\cal F}^{(1)}}_{\rm(1\ell,pow.)}(\sigma)=\frac{1}{\beta_{0}}\int_{0}^{\infty}du\;\exp\left(-\frac{u}{\beta_{0}a_{\rm(1\ell)}(\kappa\sigma)}\right)\left({\widetilde{B}}[{{\cal D}^{(1)}}](u;\kappa)\frac{\sin(\pi u)}{\pi u}\right),$ (50)
and thus
$B[{{\cal F}^{(1)}}_{\rm(1\ell,pow.)}](u;\kappa)={\widetilde{B}}[{{\cal D}^{(1)}}](u;\kappa)\frac{\sin(\pi u)}{\pi u}.$ (51)
Here we have on the left-hand side the quantity Eq. (46b), i.e., this is identical to the modified Borel ${\widetilde{B}}[{{\cal F}^{(1)}}](u;\kappa)$.
This concludes the proof of Theorem 5.
In this context, we mention that an analogous relation to Eq. (43) was obtained in BY1992 between the (usual) Borel transforms of the $R_{ee}(\sigma)$ ratio and of the Adler function in the large-$\beta_{0}$ approximation, and in BenekeDis between the Borel transforms of the $\tau$-lepton semihadronic decay ratio $R_{\tau}$ and of the Adler function in the large-$\beta_{0}$ approximation.
We recall that Theorem 5 above relates the modified Borel of the timelike and spacelike auxiliary quantities ${{\cal F}^{(1)}}(\sigma)$ Eq. (40b) and ${{\cal D}^{(1)}}(Q^{2})$ Eq. (11).
We will now prove the following analogous Theorem 6, which relates the modified Borel of the full timelike and spacelike quantities ${\cal F}(\sigma)$ Eq. (40a) and ${\cal D}(Q^{2})$ Eq. (1):
Theorem 6: Let ${\cal D}(Q^{2})$ be the spacelike observable whose expansions are written in Eqs. (1), and ${\cal F}(\sigma)$ be the corresponding timelike quantity defined in Eq. (40a), and let us assume that the modified Borel of ${\cal D}(Q^{2})$ has the form
${\widetilde{B}}^{(\nu_{0})}[{\cal D}](u;\kappa)\left[=\sum_{n=0}^{\infty}\frac{{\widetilde{d}}_{n}(\nu_{0};\kappa)}{n!\beta_{0}^{n}}u^{n}\right]$ $=$ $\frac{{\widetilde{\cal K}}(\kappa)}{(p-u)^{{\widetilde{s}}_{0}}}.$ (52)
Then the modified Borel of ${\cal F}(\sigma)$ has the following form:
${\widetilde{B}}^{(\nu_{0})}[{\cal F}](u;\kappa)\left[=\sum_{n=0}^{\infty}\frac{{\widetilde{f}}_{n}(\nu_{0};\kappa)}{n!\beta_{0}^{n}}u^{n}\right]$ $=$ $\frac{\sin(\pi p)}{\pi p}\frac{{\widetilde{\cal K}}(\kappa)}{(p-u)^{{\widetilde{s}}_{0}}}\left[1+{\cal O}\left((p-u)\right)\right].$ (53)
In this context, we recall that the coefficients ${\widetilde{d}}_{n}(\nu_{0};\kappa)$ and ${\widetilde{f}}_{n}(\nu_{0};\kappa)$ that define the expansions of these modified Borel transforms appear in the expansions (1b) and (42a) in generalised logarithmic derivatives ${\widetilde{a}}_{\nu_{0}+n}$.
Proof:
When we apply the resummation of ${\cal D}(Q^{2})$, as given in Theorem 1, Eq. (21), to the integral transformation (40a), and exchange the order of integration over $t$ and $\phi$, we obtain
${\cal F}(\sigma)_{\rm res.}=\int_{0}^{\infty}\frac{dt}{t}F_{{{\cal D}^{(1)}}}(t){\widetilde{\mathfrak{h}}}_{\nu_{0}}(t\sigma),$ (54)
where ${\widetilde{\mathfrak{h}}}_{\nu}(\sigma)$ is the timelike analog of the spacelike (generalised logarithmic derivative) ${\widetilde{a}}_{\nu}(Q^{2})$, defined and investigated in GCAK
${\widetilde{\mathfrak{h}}}_{\nu}(\kappa\sigma)\equiv\frac{1}{2\pi}\int_{-\pi}^{\pi}d\phi\;{\widetilde{a}}_{\nu}(\kappa\sigma e^{i\phi}).$ (55)
As shown in GCAK , this timelike coupling obeys the following differential recursion relation [completely analogous to that of ${\widetilde{a}}_{\nu}$ Eq. (5)]:
$\frac{d}{d\ln\kappa}{\widetilde{\mathfrak{h}}}_{\nu}(\kappa\sigma)=(-\beta_{0})\nu{\widetilde{\mathfrak{h}}}_{\nu+1}(\kappa\sigma).$ (56)
This allows us to Taylor-expand ${\widetilde{\mathfrak{h}}}_{\nu_{0}}(t\sigma)$ appearing in the integral in Eq. (54) around $\kappa\sigma$ [analogous to Eq. (22)]
${\widetilde{\mathfrak{h}}}_{\nu_{0}}(tQ^{2})=\sum_{n=0}^{\infty}(-\beta_{0})^{n}\ln^{n}\left(\frac{t}{\kappa}\right)\frac{\Gamma(\nu_{0}+n)}{\Gamma(\nu_{0})\Gamma(n+1)}{\widetilde{\mathfrak{h}}}_{\nu_{0}+n}(\kappa Q^{2}).$ (57)
When we insert this expansion in the integral (54), exchange the order of integration and summation, and use the sum rules Eq. (23a) to express the obtained integrals through the coefficinets ${\widetilde{d}}_{n}(\nu_{0};\kappa)$, we obtain the following expansion for ${\cal F}(\sigma)$:
${\cal F}(\sigma)=\sum_{n=0}^{\infty}{\widetilde{d}}_{n}(\nu_{0};\kappa){\widetilde{\mathfrak{h}}}_{\nu_{0}+n}(\kappa\sigma),$ (58)
which is completely analogous to the expansion (1b) for ${\cal D}(Q^{2})$, with the only difference being that the spacelike couplings ${\widetilde{a}}_{\nu_{0}+n}(\kappa Q^{2})$ is now replaced by the timelike couplings ${\widetilde{\mathfrak{h}}}_{\nu_{0}+n}(\kappa\sigma)$.
We now want to express in the expansion (58) the couplings ${\widetilde{\mathfrak{h}}}_{\nu_{0}+n}(\kappa\sigma)$ in terms of the couplings ${\widetilde{a}}_{\nu_{0}+m}(\kappa\sigma)$, in order to relate the coefficients ${\widetilde{d}}_{n}(\nu_{0};\kappa)$ appearing in Eq. (58) with the coefficients ${\widetilde{f}}_{n}(\nu_{0};\kappa)$ appearing in Eq. (42a), since the latter appear in the definition (expansion) of the sought modified Borel ${\widetilde{B}}^{(\nu_{0})}[{\cal F}](u;\kappa)$ Eq. (53). This can be obtained, when we use in the definition (55) of ${\widetilde{\mathfrak{h}}}_{\nu}(\kappa\sigma)$, for the coupling ${\widetilde{a}}_{\nu}(\kappa\sigma e^{i\phi})$ the Taylor-expansion around $\kappa\sigma$ [cf. the expansion Eq. (22)]
${\widetilde{a}}_{\nu}(\kappa\sigma e^{i\phi})=\sum_{m=0}^{\infty}(-\beta_{0})^{m}(i\phi)^{m}\frac{\Gamma(\nu+m)}{\Gamma(\nu)\Gamma(m+1)}{\widetilde{a}}_{\nu+m}(\kappa\sigma).$ (59)
We insert this expansion in the integral Eq. (55), exchange the order of integration and summation, and obtain the relation we sought
${\widetilde{\mathfrak{h}}}_{\nu}(\kappa\sigma)\equiv\sum_{m=0}^{\infty}(-\beta_{0})^{m}\frac{\Gamma(\nu+m)}{\Gamma(\nu)\Gamma(m+1)}I_{m}{\widetilde{a}}_{\nu+m}(\kappa\sigma),$ (60)
where the integrals $I_{m}$ are
$I_{m}\equiv\frac{1}{2\pi}\int_{-\pi}^{+\pi}d\phi\;(i\phi)^{m}.$ (61)
These integrals are nonzero only for even $m=2r$
$I_{2r}=(-1)^{r}\frac{\pi^{2r}}{(2r+1)}\qquad(I_{2r+1}=0).$ (62)
The relation (60) and the expressions (62) imply
${\widetilde{\mathfrak{h}}}_{\nu}(\kappa\sigma)\equiv\sum_{r=0}^{\infty}(-\beta_{0}^{2}\pi^{2})^{r}\frac{1}{(2r+1)}\frac{\Gamma(\nu+2r)}{\Gamma(\nu)\Gamma(2r+1)}{\widetilde{a}}_{\nu+2r}(\kappa\sigma).$ (63)
We now use this formula in the expansion (58) of ${\cal F}(\sigma)$, resulting in the following expansion of ${\cal F}(\sigma)$ in terms of ${\widetilde{a}}_{\nu_{0}+n}(\kappa\sigma)$:
${\cal F}(\sigma)$ $=$ $\sum_{n\geq 0}\sum_{r\geq 0}{\widetilde{d}}_{n}(\nu_{0};\kappa)(-\beta_{0}^{2}\pi^{2})^{r}\frac{\Gamma(\nu_{0}+n+2r)}{\Gamma(\nu_{0}+n)(2r+1)!}{\widetilde{a}}_{\nu_{0}+n+2r}(\kappa\sigma)$ (64a)
$=$ $\sum_{N\geq 0}{\widetilde{a}}_{\nu_{0}+N}(\kappa\sigma)\left[\sum_{r=0}^{[N/2]}(-\beta_{0}^{2}\pi^{2})^{r}\frac{\Gamma(\nu_{0}+N)}{\Gamma(\nu_{0}+N-2r)(2r+1)!}{\widetilde{d}}_{N-2r}(\nu_{0};\kappa)\right].$ (64b)
In the second identity (64b) we used the notation n=N2rn=N-2r ($0,1,\ldots$), and $[N/2]$ is the integer part of $N/2$. The result (64b), in conjunction with the expansion (42a), then finally gives us the sought relation between the ${\widetilde{f}}_{n}$ and ${\widetilde{d}}_{n}$ coefficients
${\widetilde{f}}_{n}(\nu_{0};\kappa)=\sum_{r=0}^{[n/2]}(-\beta_{0}^{2}\pi^{2})^{r}\frac{\Gamma(\nu_{0}+n)}{\Gamma(\nu_{0}+n-2r)(2r+1)!}{\widetilde{d}}_{n-2r}(\nu_{0};\kappa).$ (65)
Since ${\widetilde{B}}^{(\nu_{0})}[{\cal D}](u;\kappa)$ has the the form (52), this implies that ${\widetilde{d}}_{n}(\nu_{0};\kappa)$ has the asymptotic form as given in Eq. (32). We use this form in the relations (65) in order to obtain the asymptotic form for the coefficients ${\widetilde{f}}_{n}(\nu_{0};\kappa)$
${\widetilde{f}}_{n}(\nu_{0};\kappa)$ $=$ $\frac{{\widetilde{\cal K}}(\kappa)p^{-{\widetilde{s}}_{0}}}{\Gamma({\widetilde{s}}_{0})}\sum_{r=0}^{[n/2]}(-\beta_{0}^{2}\pi^{2})^{r}\left(\frac{\beta_{0}}{p}\right)^{n-2r}\frac{(n-2r)^{{\widetilde{s}}_{0}-1}}{(2r+1)!}(n-2r)!(\nu_{0}+n-1)\cdots(\nu_{0}+n-2r)\left[1+{\cal O}\left(\frac{1}{(n-2r)}\right)\right]$ (66a)
$=$ $\frac{{\widetilde{\cal K}}(\kappa)p^{-{\widetilde{s}}_{0}}}{\Gamma({\widetilde{s}}_{0})}\left(\frac{\beta_{0}}{p}\right)^{n}n!n^{{\widetilde{s}}_{0}-1}{\Bigg{\{}}1+\ldots+\frac{(-\beta_{0}^{2}\pi^{2})^{r}}{(2r+1)!}\left(\frac{p}{\beta_{0}}\right)^{2r}\left(1-\frac{(1-\nu_{0})}{n}\right)\cdots\left(1-\frac{(1-\nu_{0})}{(n-2r+1)}\right)$
$\times\left[1+{\cal O}\left(\frac{1}{(n-2r)}\right)\right]+\ldots{\Bigg{\}}}$
$=$ $\frac{{\widetilde{\cal K}}(\kappa)p^{-{\widetilde{s}}_{0}}}{\Gamma({\widetilde{s}}_{0})}\left(\frac{\beta_{0}}{p}\right)^{n}n!n^{{\widetilde{s}}_{0}-1}\left\{1+\ldots+\frac{(-\beta_{0}^{2}\pi^{2})^{r}}{(2r+1)!}\left(\frac{p}{\beta_{0}}\right)^{2r}\left[1+{\cal O}\left(\frac{1}{(n-2r)}\right)+{\cal O}\left(\frac{r}{(n-2r)}\right)\right]+\ldots\right\}$ (66b)
$\approx$ $\frac{{\widetilde{\cal K}}(\kappa)p^{-{\widetilde{s}}_{0}}}{\Gamma({\widetilde{s}}_{0})}\left(\frac{\beta_{0}}{p}\right)^{n}n!n^{{\widetilde{s}}_{0}-1}\left\{1+\ldots+\frac{(-p\pi)^{2r}}{(2r+1)!}+\ldots\right\}\left[1+{\cal O}\left(\frac{1}{n}\right)\right]$ (66c)
$=$ $\frac{{\widetilde{\cal K}}(\kappa)p^{-{\widetilde{s}}_{0}}}{\Gamma({\widetilde{s}}_{0})}\left(\frac{\beta_{0}}{p}\right)^{n}n!n^{{\widetilde{s}}_{0}-1}\frac{\sin(\pi p)}{\pi p}\left[1+{\cal O}\left(\frac{1}{n}\right)\right].$ (66d)
In Eq. (66c) we took into account that $n\gg 1$, and that in the sum over $r$ only terms $r\ll n$ contribute, because of the strong suppression of the terms by the factor $1/(2r+1)!$. In fact, as we see in Eqs. (66c)-(66d), the sum over $r$ converges fast to $\sin(\pi p)/(\pi p)$. When we compare this result with the asymptotic behaviour of ${\widetilde{d}}_{n}(\nu_{0};\kappa)$ Eq. (32), we see that the result (66d) implies the relation
${\widetilde{f}}_{n}(\nu_{0};\kappa)=\frac{\sin(\pi p)}{\pi p}{\widetilde{d}}_{n}(\nu_{0};\kappa)\left(1+{\cal O}(1/n)\right),$ (67)
and this then immediately implies Eq. (53). This concludes the proof of Theorem 6.
One of the consequences of Theorems 5 and 6 is the following lemma:
[…]
${\widetilde{f}}_{n}(1;\kappa)=\frac{\Gamma(\nu_{0})\Gamma(1+n)}{\Gamma(\nu_{0}+n)}{\widetilde{f}}_{n}(\nu_{0};\kappa)\left(1+{\cal O}(1/n)\right),$ (68)
which is asymptotically analogous to the relations (9) of the $d_{n}$ coefficients of the corresponding spacelike quantities $\cal D$} and ${\cal D}^{(1)}$}.
Proof:
The form Eq. (52) for ${\widetilde{B}}^{(\nu_{0})}[{\cal D}](u;\kappa)$ implies, according to Theorem 3, Eq. (28), a specific form for ${\widetilde{B}}[{{\cal D}^{(1)}}](u;\kappa)$, namely
${\widetilde{B}}^{(\nu_{0})}[{\cal D}](u;\kappa)$ $=$ $\frac{{\widetilde{\cal K}}(\kappa)}{(p-u)^{{\widetilde{s}}_{0}}}\left[1+{\cal O}((p-u))\right],\;\Rightarrow$ (69a)
${\widetilde{B}}[{{\cal D}^{(1)}}](u;\kappa)$ $=$ $\frac{{{\widetilde{\cal K}}^{(1)}}(\kappa)}{(p-u)^{{\widetilde{s}}_{0}-\nu_{0}+1}}\left[1+{\cal O}((p-u))\right].$ (69b)
In conjunction with Theorem 5, Eq. (43), this implies
${\widetilde{B}}[{{\cal F}^{(1)}}](u;\kappa)$ $=$ $\frac{\sin(\pi u)}{\pi u}\frac{{{\widetilde{\cal K}}^{(1)}}(\kappa)}{(p-u)^{{\widetilde{s}}_{0}-\nu_{0}+1}}\left[1+{\cal O}((p-u))\right]$ (70a)
$=$ $\frac{\sin(\pi p)}{\pi p}\frac{{{\widetilde{\cal K}}^{(1)}}(\kappa)}{(p-u)^{{\widetilde{s}}_{0}-\nu_{0}+1}}\left[1+{\cal O}((p-u))\right],$ (70b)
where in the last identity we took into account that
$\frac{\sin(\pi u)}{\pi u}=\frac{\sin(\pi p)}{\pi p}\left(1+{\cal O}(p-u)\right).$ (71)
The relation (70b), in conjunction with the notations (11) and (42b) for the expansions of ${\cal D}^{(1)}$} and ${\cal F}^{(1)}$}, respectively, then implies
${\widetilde{f}}_{n}(1;\kappa)=\frac{\sin(\pi p)}{\pi p}{\widetilde{d}}_{n}(1;\kappa)\left(1+{\cal O}(1/n)\right).$ (72)
This relation, together with the analogous relation (67) for ${\widetilde{f}}_{n}(\nu_{0};\kappa)$ and ${\widetilde{d}}_{n}(\nu_{0};\kappa)$ [that reflects Theorem 6, Eqs. (52)-(53)], and the original relation (9) between ${\widetilde{d}}_{n}(1;\kappa)$ and ${\widetilde{d}}_{n}(\nu_{0};\kappa)$, then immediately implies the relation (68), i.e., the claim of the Lemma. This concludes the proof of the Lemma.
IV The use of holomorphic QCD (AQCD) couplings in the formalism
The holomorphic (AQCD) coupling ${\mathcal{A}}(Q^{2})$ [i.e., when $a(Q^{2})$ is replaced by ${\mathcal{A}}(Q^{2})$] is such that ${\mathcal{A}}(Q^{2})$ is a holomorphic (analytic) function of $Q^{2}$ in the entire complex $Q^{2}$-plane with the exception of the negative semiaxis, i.e., for all $Q^{2}\in\mathbb{C}\backslash(-\infty,-M^{2}_{\rm thr})$ where $M_{\rm thr}\sim m_{\pi}\sim 0.1\ {\rm GeV}$ is a threshold mass. Such a behaviour qualitatively reflects the holomorphic properties of the (QCD) spacelike observables ${\cal D}(Q^{2})$, where the latter properties are a direct consequence of the locality, unitarity and causality of Quantum Field Theories BookShirkov . This is in contrast to the properties of the pQCD coupling $a(Q^{2})$ which has (Landau) cuts within the Euclidean regime of $Q^{2}$, usually on the positive axis: $0\leq Q^{2}\leq\Lambda^{2}_{\rm Lan}$; $\Lambda_{\rm Lan}\sim 0.1\ {\rm GeV}$. A basic property of ${\mathcal{A}}(Q^{2})$ is that it should effectively coincide with the perturbative $a(Q^{2})$ for large Euclidean $Q^{2}$ ($Q^{2}>1{\rm GeV}^{2}$).
Usually the AQCD coupling framework is defined via the specification of the form of the spectral (discontinuity) function of the coupling $\mathcal{A}$} along its cut: $\rho_{{\mathcal{A}}}(\sigma)\equiv{\rm Im}\;{\mathcal{A}}(-\sigma-i\varepsilon)$ for all $\sigma\geq M^{2}_{\rm thr}$. At large squared timelike momenta $\sigma$, this function should effectively coincide with its pQCD counterpart, $\rho_{{\mathcal{A}}}(\sigma)=\rho_{a}(\sigma)$. Here, $a(Q^{2})$ is called the underlying pQCD coupling; $a(Q^{2})$ and ${\mathcal{A}}(Q^{2})$ are in the same renormalisation scheme.
This approach suggests that the construction of an AQCD framework reduces basically to the modelling of the spectral function $\rho_{{\mathcal{A}}}(\sigma)$ at low $\sigma$: $0<\sigma<1\ {\rm GeV}^{2}$. [^16] The most prominent early model of holomorphic coupling has the spectral function $\rho_{{\mathcal{A}}}(\sigma)$ equal to the pQCD $\rho_{a}(\sigma)$ for all $\sigma>0$, including low values $\sigma\lesssim 1\ {\rm GeV}^{2}$. This model can be called the Minimal Analytic framework (MA; named also (F)APT) Shirkov ; SMS ; ShirRev ; FAPT ; BakRev , and thus has $\rho_{{\mathcal{A}}}^{\rm(MA)}(\sigma)=\Theta(\sigma)\rho^{\rm(pt)}(\sigma)$. The difference ${\mathcal{A}}(Q^{2})-a(Q^{2})$ in the MA at large $|Q^{2}|$ is $\sim(\Lambda^{2}_{\rm Lan}/Q^{2})^{1}$, i.e., a relatively slow fall-off when $|Q^{2}|$ increases. A generalisation of (F)APT was proposed in AM1 ; AM2 in the context of the pion form factor. In this regime, it is reasonable to represent the parametrisation of $\rho_{{\mathcal{A}}}(\sigma)$ in terms of a sum of a finite number of Dirac deltas, e.g. two deltas 2dQCD1 ; 2dQCD2 or three deltas 3dAQCD .
[…]
$\rho_{{\mathcal{A}}}(\sigma)=\pi\sum_{j=1}^{3}{{\cal F}}_{j}\delta(\sigma-M_{j}^{2})+\Theta(\sigma-M_{0}^{2})\rho_{a}(\sigma),$ (73)
($\Theta$ being the Heaviside step function) and we assume the following hierarchy of the squared masses: $0<M_{1}^{2}<M_{2}^{2}<M_{3}^{2}<M_{0}^{2}$. Here, $M_{1}^{2}=M^{2}_{\rm thr}$ is the IR-threshold scale (i.e., $\sigma_{\rm min}=M_{1}^{2}$), and $M_{0}^{2}$ is the pQCD-onset scale ($M_{0}\sim 1$ GeV). The holomorphic coupling is then obtained, with the use of the Cauchy theorem, as a dispersion integral involving $\rho_{{\mathcal{A}}}(\sigma)$
${\mathcal{A}}(Q^{2})$ $=$ $\frac{1}{\pi}\int_{-M^{2}_{\rm thr}-\eta}^{\infty}\frac{d\sigma\;\rho_{{\mathcal{A}}}(\sigma)}{(\sigma+Q^{2})}\qquad(\eta\to+0)$ (74a)
$=$ $\sum_{j=1}^{3}\frac{{{\cal F}}_{j}}{(Q^{2}+M_{j}^{2})}+\frac{1}{\pi}\int_{M^{2}_{0}}^{\infty}\frac{d\sigma\;\rho_{a}(\sigma)}{(\sigma+Q^{2})}.$ (74b)
The pQCD coupling $a(Q^{2})$ (in the LMM scheme) has $n_{f}=3$ or $n_{f}=4$, if we are interested in low-energy QCD phenomenology. This $a(Q^{2})$, and thus also its spectral function $\rho_{a}(\sigma)\equiv{\rm Im}\;a(-\sigma-i\varepsilon)$, is completely fixed by specifying the value of $\alpha_{s}^{(\overline{\rm MS})}(M_{Z}^{2})$.
For specifying numerically the other seven parameters (${\cal F}_{j}$, $M_{j}^{2}$ for $j=1,2,3$; and $M_{0}^{2}$), we use the following inputs: For large $Q^{2}$ ($|Q^{2}|>1\ {\rm GeV}^{2}$) we require ${\mathcal{A}}(Q^{2})$ to approach the underlying perturbative coupling $a(Q^{2})$ quickly, specifically ${\mathcal{A}}(Q^{2})-a(Q^{2})\sim(\Lambda_{\rm Lan}^{2}/Q^{2})^{5}$, and this gives us four conditions; whereas at low $Q^{2}$ the required behaviour is as suggested by large-volume lattice calculations 3dLVLattice : ${\mathcal{A}}(Q^{2})$ at positive $Q^{2}$ has a local maximum at $Q^{2}\approx 0.135\ {\rm GeV}^{2}$ and for $Q^{2}\to 0$ it behaves as ${\mathcal{A}}(Q^{2})\sim Q^{2}$ ($\to 0$). One of the seven parameters is then still free, which we choose to be the threshold scale of the spectral function, namely $\sigma_{\rm thr.}=M_{1}^{2}$, which is expected to be of the order of the square of the lowest hadronic mass: $M_{1}^{2}\approx m_{\pi}^{2}$ ($\sim 0.15^{2}\ {\rm GeV}^{2}$). Note that the renormalisation scheme we use in this construction is MiniMOM MiniMOM because the large-volume lattice calculations (3dLVLattice ) are performed in this scheme. In addition, we rescale the momenta to the usual $\overline{\rm MS}$-type scheme scales (scaled with $\Lambda^{2}_{\overline{\rm MS}}$), and this is called Lambert MiniMOM (LMM) scheme. We refer for details of construction of ${\mathcal{A}}(Q^{2})$ in such framework to 3dAQCD ; amu3dAQCD .
In AQCD, the generalised logarithmic derivatives in pQCD, ${\widetilde{a}}_{\nu}(Q^{2})$ [cf. Eq. (2a)], get replaced by their AQCD analogs ${\widetilde{\mathcal{A}}}_{\nu}(Q^{2})$ whose expression turns out to be, conveniently, determined entirely by a dispersive integral involving the spectral function $\rho_{{\mathcal{A}}}(\sigma)$ GCAK  [^17] The use of the logarithmic derivatives ${\widetilde{\mathcal{A}}}_{\nu}$ for integer $\nu=n$ for any AQCD framework was developed in CV , and was extended to noninteger $\nu$ in GCAK (for any AQCD). For the Minimal Analytic (MA) QCD, the extended logarithmic derivatives ${\widetilde{\mathcal{A}}}_{\nu}^{\rm(MA)}(Q^{2})$ were constructed, in a MA-specific way, as explicit functions at one-loop order in FAPT and at any loop order in Kotikov . We point out that the general approach GCAK , Eq. (75), can also be applied in the MA case and it gives the same numerical result as the MA-specific approach.
${\widetilde{\mathcal{A}}}_{\nu}(Q^{2})$ $=$ $\frac{1}{\pi}\frac{(-1)}{\beta_{0}^{\nu-1}\Gamma(\nu)}\int_{0}^{\infty}\frac{d\sigma}{\sigma}\rho_{{\mathcal{A}}}(\sigma){\rm Li}_{-\nu+1}\left(-\frac{\sigma}{Q^{2}}\right)\quad(\nu>0).$ (75)
This formula can even be modified, by a subtraction approach GCAK , so that is becomes valid for even lower values of the index $\nu$ ($\nu>-1$).
Furthermore, the timelike analog of this coupling, namely ${\widetilde{\mathfrak{H}}}(\sigma)$ which is the AQCD analog of ${\widetilde{\mathfrak{h}}}_{\nu}$ of Eq. (55), can also be obtained in terms of an integral involving the spectral function $\rho_{{\mathcal{A}}}$ GCAK
${\widetilde{\mathfrak{H}}}_{\nu}(\kappa\sigma)$ $\equiv$ $\frac{1}{2\pi}\int_{-\pi}^{\pi}d\phi\;{\widetilde{\mathcal{A}}}_{\nu}(\kappa\sigma e^{i\phi}).$ (76a)
$=$ $-\frac{\sin(\pi\nu)}{\pi^{2}(\nu-1)\beta_{0}^{\nu-1}}\int_{0}^{\infty}\frac{dw}{w^{\nu-1}}\rho_{{\mathcal{A}}}(\sigma e^{w})\quad(0<\nu<2),$ (76b)
[…]
${\widetilde{\mathfrak{H}}}_{\nu}(\kappa\sigma)=\frac{\sin(\pi\nu)\beta_{0}^{1-\nu}}{\pi^{2}(1-\nu)}\left\{\pi\sum_{j=1}^{3}\frac{{\cal F}_{j}}{M_{j}^{2}}\Theta(M_{j}^{2}-\sigma)\ln^{1-\nu_{0}}\left(\frac{M_{j}^{2}}{\sigma}\right)+\int_{\Theta(M_{0}^{2}-\sigma)\ln(M_{0}^{2}/\sigma)}^{\infty}dw\;w^{1-\nu_{0}}\rho_{a}(\sigma e^{w})\right\}.$ (77)
All the results of the previous Sections with pQCD couplings can now be simply rewritten in AQCD, by the replacements: [^18] We point out, though, that these results are evaluated in practice in (any) AQCD more simply than in pQCD, because in AQCD the basic elements are $\mathcal{A}$} and the generalised logarithmic derivatives ${\widetilde{\mathcal{A}}}_{\nu}$ and ${\widetilde{\mathfrak{H}}}_{\nu}$ (determined entirely by $\rho_{{\mathcal{A}}}(\sigma)$), while this is not true for pQCD (where the powers of $a$ are the basic elements). $a\mapsto{\mathcal{A}}$, ${\widetilde{a}}_{\nu}\mapsto{\widetilde{\mathcal{A}}}_{\nu}$, and ${\widetilde{\mathfrak{h}}}_{\nu}\mapsto{\widetilde{\mathfrak{H}}}_{\nu}$. The central results for renormalon-motivated resummations, Eqs. (39), are now simply rewritten for (any) AQCD in the following form:
${{\cal D}^{(1)}}(Q^{2})_{\rm res.}$ $=$ $\int_{0}^{\infty}\frac{dt}{t}F_{{{\cal D}^{(1)}}}(t)_{(p,{\widetilde{s}})}{\mathcal{A}}(te^{-{\widetilde{K}}_{e}}Q^{2}),$ (78a)
${\cal D}(Q^{2})_{\rm res.}$ $=$ $\int_{0}^{\infty}\frac{dt}{t}F_{{{\cal D}^{(1)}}}(t)_{(p,{\widetilde{s}})}{\widetilde{\mathcal{A}}}_{\nu_{0}}(te^{-{\widetilde{K}}_{e}}Q^{2}),$ (78b)
${\cal F}(\sigma)_{\rm res.}$ $=$ $\int_{0}^{\infty}\frac{dt}{t}F_{{{\cal D}^{(1)}}}(t)_{(p,{\widetilde{s}})}{\widetilde{\mathfrak{H}}}_{\nu_{0}}(te^{-{\widetilde{K}}_{e}}\sigma).$ (78c)
Furthermore, due to the Landau cuts of the pQCD coupling $a(Q^{2})$ at low positive $Q^{2}$ ($0\leq Q^{2}\leq\Lambda^{2}_{\rm Lan}$) we have to regularise the pQCD resummations Eqs. (39) is some way to avoid those cuts in the integration of $t$ at low $t$ values. One possibility is a PV-type regularisation
${{\cal D}^{(1)}}(Q^{2})_{\rm res.}$ $=$ ${\rm Re}\;\int_{0}^{\infty}\frac{dt}{t}F_{{{\cal D}^{(1)}}}(t)_{(p,{\widetilde{s}})}a(te^{-{\widetilde{K}}_{e}}Q^{2}+i\varepsilon),$ (79a)
${\cal D}(Q^{2})_{\rm res.}$ $=$ ${\rm Re}\;\int_{0}^{\infty}\frac{dt}{t}F_{{{\cal D}^{(1)}}}(t)_{(p,{\widetilde{s}})}{\widetilde{a}}_{\nu_{0}}(te^{-{\widetilde{K}}_{e}}Q^{2}+i\varepsilon),$ (79b)
${\cal F}(\sigma)_{\rm res.}$ $=$ ${\rm Re}\;\int_{0}^{\infty}\frac{dt}{t}F_{{{\cal D}^{(1)}}}(t)_{(p,{\widetilde{s}})}{\widetilde{\mathfrak{h}}}_{\nu_{0}}(te^{-{\widetilde{K}}_{e}}\sigma+i\varepsilon),$ (79c)
where $\varepsilon\to+0$.
In pQCD, one may argue that for the resummation of a timelike quantity ${\cal F}(\sigma)$ we may use exactly the same approach as for a spacelike quantity ${\cal D}(Q^{2})$: namely, if the perturbation expansions of these quantities, one in powers of $a(\sigma)$ and the other in powers of $a(Q^{2})$, have the same structure, cf. Eqs. (1) and (42a), [^19] Eq. (42a) is a series in generalised logarithmic derivatives ${\widetilde{a}}_{\nu_{0}+n}(\kappa\sigma)$, but it can be rewritten in powers, namely as $\sum f_{n}(\nu_{0};\kappa)a(\kappa\sigma)^{\nu_{0}+n}$, in exactly the same way that ${\cal D}(Q^{2})$ was written in powers of $a(\kappa Q^{2})^{\nu_{0}+n}$ in Eq. (1a). then we can construct the auxiliary quantity ${\cal F}^{(1)}(\sigma)$ and the corresponding characteristic function $F_{{\cal F}^{(1)}}(t)$ that would then appear in the integral of $t$ involving also the factor ${\widetilde{a}}_{\nu_{0}}(te^{-{\widetilde{K}}_{e}}\sigma)$ $\mapsto{\widetilde{\mathcal{A}}}_{\nu_{0}}(te^{-{\widetilde{K}}_{e}}\sigma)$]. However, the work GCAK makes it clear that the couplings $a$ and ${\widetilde{a}}_{\nu_{0}}$ have their AQCD analogs $\mathcal{A}$}, ${\widetilde{\mathcal{A}}}_{\nu_{0}}$ that are holomorphic functions of $Q^{2}$ in the complex-$Q^{2}$ plane (with the exception of the negative semiaxis), and that they reflect in this way the holomorphic properties of the spacelike QCD observables. The timelike QCD observables ${\cal F}(\sigma)$ have no such holomorphic properties, they are defined only for $\sigma>0$, and they or their derivatives are in general not even continuous functions of $\sigma$. Therefore, the use of $\mathcal{A}$} and ${\widetilde{\mathcal{A}}}_{\nu_{0}}$ (in pQCD: of $a$ and ${\widetilde{a}}_{\nu_{0}}$) couplings for ${\cal F}(\sigma)$ is not warranted.
[…]
When we change the renormalisation scheme (i.e., we change the scheme parameters $c_{j}$, $j\geq 2$), e.g., from $\overline{\rm MS}$ to LMM, the expansion coefficients $f_{n}$ (and $d_{n}$) transform according to the rules as given in Appendix C. We point out that the LMM scheme involves, in our convention, no rescaling of the momenta (i.e., it remains in that sense a $\overline{\rm MS}$-type scheme), only the scheme parameters $c_{2}$ and $c_{3}$ change.
[…]
$M_{1}$ [GeV] $\alpha_{s}^{\overline{\rm MS}}(M_{Z}^{2})$ $\Lambda_{{\rm L}}$ [GeV] $f_{1}$ $f_{2}$ $f_{3}$ $s_{1}$ $s_{2}$ $s_{3}$ $s_{0}$
0.100 0.1180 0.112500 -0.168833 12.21307 7.50079 0.790117 87.4086 858.412 1158.76
0.150 0.1180 0.112500 -0.583466 10.64786 6.055440 1.77776 42.6800 605.184 824.850
0.250 0.1180 0.112500 -4.35108 13.24663 5.22514 4.93823 16.58907 469.278 645.569
0.150 0.1171 0.108036 -0.609272 10.82440 6.16694 1.92773 45.5255 623.641 849.268
0.150 0.1189 0.117067 -0.559701 10.48340 5.95174 1.64178 40.0474 588.119 802.275
Table 2: The same as in Table 1, but now with the underlying pQCD coupling (in the LMM scheme) with $n_{f}=4$.
$M_{1}$ [GeV] $\alpha_{s}^{\overline{\rm MS}}(M_{Z}^{2})$ $\Lambda_{{\rm L}}$ [GeV] $f_{1}$ $f_{2}$ $f_{3}$ $s_{1}$ $s_{2}$ $s_{3}$ $s_{0}$
0.100 0.1180 0.0960725 -0.225523 14.44965 8.98645 1.08343 112.1859 981.749 1319.90
0.150 0.1180 0.0960725 -0.757927 12.28852 7.00460 2.43772 54.4228 659.170 894.669
0.250 0.1180 0.0960725 -5.615187 15.55630 5.86556 6.77145 21.0441 488.475 669.434
0.150 0.1171 0.0917892 -0.799402 12.56215 7.17931 2.67054 58.4764 685.697 929.797
0.150 0.1189 0.100482 -0.719985 12.03468 6.84274 2.22845 50.6968 634.799 862.397
[…]
We recall that the spacelike couplings, such as ${\widetilde{\mathcal{A}}}_{\nu_{0}}(Q^{2})$, are holomorphic functions of $Q^{2}$ in the $Q^{2}$-complex plane $Q^{2}\in\mathbb{C}\backslash(-\infty,-M^{2}_{\rm thr})$, where $M_{\rm thr}=M_{1}$ is the IR-threshold scale of the spectral function $\rho_{{\mathcal{A}}}(\sigma)$, Eq. (73). On the other hand, the timelike couplings, such as ${\widetilde{\mathfrak{H}}}_{\nu_{0}}(\sigma)$, are defined only for positive $\sigma\geq 0$, and are in general neither holomorphic nor are their derivatives continuous, especially when the low-$\sigma$ regime of $\rho_{{\mathcal{A}}}(\sigma)$ is parametrised by Dirac delta functions. In fact, the derivatives of ${\widetilde{\mathfrak{H}}}_{\nu_{0}}(\sigma)$ are discontinuous at $\sigma=M_{1}^{2},M_{2}^{2},M_{3}^{2}$ (although the discontinuity at the highest squared mass $M_{3}^{2}\sim 10\ {\rm GeV}^{2}$ is weak and cannot be seen by simple inspection).
V An example: Implementation of resummation, timelike case
Here we will show how to implement the renormalon-motivated resummation in a case of a timelike observable. The considered observable will be ${\cal F}(\sigma)={\hat{C}}(m)$ of Ref. GandN , which is the renormalisation-scheme invariant factor of the Wilson coefficient of the chromomagnetic operator in the heavy-quark effective theory (HQET) for hadronic bound states containing one heavy quark ($c$ or $b$). Strictly speaking, ${\cal F}(\sigma)=\pi^{-\nu_{0}}{\hat{C}}(m)$, where $\sigma=m$ is the (pole) mass of the heavy quark, and $\nu_{0}=\gamma_{0}/(8\beta_{0})$ is the (noninteger) power index in the expansion of ${\cal F}(\sigma)$ Eq. (42a). Here, $\gamma_{0}=2C_{A}=6$ is the one-loop coefficient of the anomalous dimension of the chromomagnetic operator, and $\beta_{0}=(11-2n_{f}/3)/4$ is the aforementioned one-loop coefficient of the $\beta$-function [cf. Eqs. (6)], where $n_{f}$ is the number of active light quark flavours. For example, in the case of $n_{f}=3$ ($m=m_{c}\approx 1.67\ {\rm GeV}$) we have $\nu_{0}=1/3$.
The expansion in (generalised) logarithmic derivatives of ${\cal F}(\sigma)$ is given in Eq. (42a). Here we write this expansion, together with the usual expansion in powers of $a$:
${\cal F}(\sigma)$ $=$ $\sum_{n=0}^{\infty}f_{n}(\nu_{0};\kappa)a(\kappa\sigma)^{\nu_{0}+n}\;\left[=\sum_{n=0}^{\infty}{\widetilde{f}}_{n}(\nu_{0};\kappa){\widetilde{a}}_{\nu_{0}+n}(\kappa\sigma)\right].$ (80)
We recall that these two expansions are completely analogous with those of the corresponding spacelike observable ${\cal D}(Q^{2})$, Eqs. (1).
It turns out that the coefficients $f_{n}(\nu_{0};\kappa)$ are known212121 In GandN the relevant coefficients are denoted as $c_{n}$, and they are related to our $f_{n}(\nu_{0};\kappa)$ ($\kappa=1$) as: $f_{n}(\nu_{0};1)=c_{n}/4^{n}$ ($n=0,1,\ldots$). Their quantity ${\hat{C}}(m_{q})$ is related to ${\cal F}(\sigma)$ as: ${\hat{C}}(m_{q})=$ $\pi^{\nu_{0}}{\cal F}(m_{q}^{2})$. for $n=0,1$ GandN :
$f_{0}(\nu_{0};\kappa)$ $=$ $1,$ (81a)
$f_{1}(\nu_{0};1)$ $=$ $\frac{1}{4}\left(\frac{91}{6}-\frac{189}{8\beta_{0}}+\frac{321}{16\beta_{0}^{2}}\right)\;\left[=\frac{233}{108}(n_{f}=3);\;\frac{7921}{3750}(n_{f}=4)\right]$ (81b)
The leading coefficient $f_{0}$ is $\kappa$-independent. The higher order coefficients $f_{2}(\nu_{0};1)$ and $f_{3}(\nu_{0};1)$ can be obtained using the results of Getal , and turn out to be
$f_{2}(\nu_{0};1)$ $=$ $16.89993447125\;(n_{f}=3);\;14.6762041125\;(n_{f}=4);$ (82a)
$f_{3}(\nu_{0};1)$ $=$ $193.419605571875\;(n_{f}=3);\;150.11891031953124\;(n_{f}=4).$ (82b)
The numerical results for $f_{3}(\nu_{0};1)$ are approximate, because in order to obtain the exact value, the value of the four-loop anomalous dimension coefficient $\gamma_{3}$ is needed (which is not known). The value of $f_{3}(\nu_{0};1)$ we give here is obtained by taking $\gamma_{3}=0$. However, there are indications that this approximation is reasonably good, as it gives in the large-$\beta_{0}$ approximation a value that is close to the known large-$\beta_{0}$ value. [^22] The large-$\beta_{0}$ value is $f_{3}^{(LB)}(\nu_{0};1)\approx 56.6608\beta_{0}^{2}$, while for the higher power on $\beta_{0}$ of the exact coefficient is $f_{3}(\nu_{0};1)\approx 57.2618\beta_{0}^{2}+\mathcal{O}(\beta_{0})$ for $\gamma_{3}=0$. We stress that the large-$\beta_{0}$ approximation is recovered by taking $\gamma_{3}=\gamma_{3}^{(LB)}$.
The Borel transform $B^{(\nu_{0})}[{\cal F}](u;\kappa)$, is defined in our convention as [cf. Eq. (24a) for Borel of ${\cal D}(Q^{2})$]
$B^{(\nu_{0})}[{\cal F}](u;\kappa)\equiv\sum_{n=0}^{\infty}\frac{f_{n}(\nu_{0};\kappa)}{n!\beta_{0}^{n}}u^{n}.$ (83)
The corresponding inverse Borel is then [cf. expansion (80)]
${\cal F}(\sigma)=a(\kappa\sigma)^{\nu_{0}-1}\frac{1}{\beta_{0}}\int_{0}^{\infty}du\exp\left(-\frac{u}{\beta_{0}a(\kappa\sigma)}\right)B^{(\nu_{0})}[{\cal F}](u;\kappa).$ (84)
The Borel transform has then, according to GandN , [^23] The Borel $S(u)$ given in Eq. (32) of GandN corresponds, in our conventions, to our Borel via the relation: $e^{5u/3}S(u)=4\beta_{0}(d/du)B^{(\nu_{0})}[{\cal F}](u;\kappa)$ with $\kappa=1$. the following leading $u=1/2$ IR renormalon terms
$B^{(\nu_{0})}[{\cal F}](u;1)=\left\{\frac{S_{+}}{\left(\frac{1}{2}-u\right)^{+\nu_{0}+\beta_{1}/(2\beta_{0}^{2})}}+\frac{S_{0}}{\left(\frac{1}{2}-u\right)^{+\beta_{1}/(2\beta_{0}^{2})}}+\frac{S_{-}}{\left(\frac{1}{2}-u\right)^{-\nu_{0}+\beta_{1}/(2\beta_{0}^{2})}}\right\}\left[1+{\cal O}\left(\frac{1}{2}-u\right)\right].$ (85)
[…]
${\widetilde{B}}^{(\nu_{0})}[{\cal F}](u;1)=\left\{\frac{{\widetilde{S}}_{+}}{\left(\frac{1}{2}-u\right)^{\nu_{0}}}+{\widetilde{S}}_{0}\ln\left(\frac{1}{2}-u\right)+\frac{{\widetilde{S}}_{-}}{\left(\frac{1}{2}-u\right)^{-\nu_{0}}}\right\}\left[1+{\cal O}\left(\frac{1}{2}-u\right)\right].$ (86)
We note that the logarithmic term appears because it corresponds effectively to the zero power renormalon term
$\ln\left(\frac{1}{2}-u\right)=\lim\limits_{\varepsilon\to+0}(-\frac{1}{\varepsilon})\left[\frac{1}{\left(\frac{1}{2}-u\right)^{\varepsilon}}-1\right].$ (87)
[…]
${\widetilde{B}}[{{\cal F}^{(1)}}](u;\kappa)=\left\{\frac{{\widetilde{S}}_{+}^{(1)}}{\left(\frac{1}{2}-u\right)^{1}}+\frac{{\widetilde{S}}_{0}^{(1)}}{\left(\frac{1}{2}-u\right)^{-\nu_{0}+1}}+\frac{{\widetilde{S}}_{-}^{(1)}}{\left(\frac{1}{2}-u\right)^{-2\nu_{0}+1}}\right\}\left[1+{\cal O}\left(\frac{1}{2}-u\right)\right].$ (88)
When we now apply Theorem 5, Eq. (43), where we take into account the expansion (71) of the proportionality factor, we obtain
${\widetilde{B}}[{{\cal D}^{(1)}}](u;\kappa)=\left\{\frac{{\widetilde{K}}_{+}^{(1)}(\kappa)}{\left(\frac{1}{2}-u\right)^{1}}+\frac{{\widetilde{K}}_{+0}^{(1)}(\kappa)}{\left(\frac{1}{2}-u\right)^{-\nu_{0}+1}}+\frac{{\widetilde{K}}_{-}^{(1)}(\kappa)}{\left(\frac{1}{2}-u\right)^{-2\nu_{0}+1}}\right\}\left[1+{\cal O}\left(\frac{1}{2}-u\right)\right],$ (89)
where ${\widetilde{K}}^{(1)}_{q}={\widetilde{S}}^{(1)}_{q}(\pi p)/(\sin(\pi p))$. When we take into account the dependence Eq. (14) under the variation of the renormalisation scale parameter $\kappa\equiv\mu^{2}/Q^{2}$, this implies the following $\kappa$-dependence of these parameters:
${\widetilde{K}}_{q}^{(1)}(\kappa)=\exp\left(\ln(\kappa)u\right){\widetilde{K}}_{q}^{(1)}\quad(q=+,0,-).$ (90)
This then means that at $\kappa=1$ the modified Borel ${\widetilde{B}}[{{\cal D}^{(1)}}](u)$ in Eq. (89) has three parameters, ${\widetilde{K}}_{q}^{(1)}\equiv{\widetilde{K}}_{q}^{(1)}(1)$ ($q=+,0,$-). However, even at $\kappa=1$, we may want to allow for the dependence of this modified Borel on the redefinition (rescaling) of the momentum, which would give us the following form of ${\widetilde{B}}[{{\cal D}^{(1)}}]$ (at $\kappa=1$):
${\widetilde{B}}[{{\cal D}^{(1)}}](u)=\exp({\widetilde{K}}_{e}^{(1)}u)\left\{\frac{{\widetilde{K}}_{+}^{(1)}}{\left(\frac{1}{2}-u\right)^{1}}+\frac{{\widetilde{K}}_{0}^{(1)}}{\left(\frac{1}{2}-u\right)^{-\nu_{0}+1}}+\frac{{\widetilde{K}}_{-}^{(1)}}{\left(\frac{1}{2}-u\right)^{-2\nu_{0}+1}}\right\}.$ (91)
As indicated, the possible relative subleading corrections ${\cal O}(1/2-u)$ to this ansatz Eq. (91) will be neglected. This expression has four parameters: ${\widetilde{K}}_{q}^{(1)}$ (q=e,+,0,q=e,+,0,-).262626 The exponential factor $\exp({\widetilde{K}}_{e}^{(1)}u)$ in Eq. (91) can be interpreted as $\exp({\widetilde{K}}_{e}^{(1)}u)=\exp({\widetilde{K}}_{e}^{(1)}/2)(1-{\widetilde{K}}_{e}^{(1)}(1/2-u))$ ${\cal O}((1/2-u)^{2})$. This means that this factor generates next-to-leading relative corrections ${\cal O}(1/2-u)$ to each of the three leading renormalon terms of Eq. (89) in the way that rescaling of the momentum would generate. These four parameters can be determined by the knowledge of the original first four coefficients $f_{n}(\nu_{0};\kappa)$ (at $\kappa=1$), Eqs. (81)-(82).
[…]
$F_{{{\cal D}^{(1)}}}(t)=\Theta(1-t)t^{1/2}\left\{{\widetilde{K}}_{+}^{(1)}+\frac{{\widetilde{K}}_{0}^{(1)}}{\Gamma(-\nu_{0}+1)(-\ln t)^{\nu_{0}}}+\frac{{\widetilde{K}}_{-}^{(1)}}{\Gamma(-2\nu_{0}+1)(-\ln t)^{2\nu_{0}}}\right\}.$ (92)
The resummations are performed according to the formulas (79) in pQCD and (78) in (holomorphic) AQCD, where $F_{{{\cal D}^{(1)}}}(t)_{(p,{\widetilde{s}})}$ is replaced by $F_{{{\cal D}^{(1)}}}(t)$ of Eq. (92). Specifically, the sought resummation of the considered timelike quantity ${\cal F}(\sigma)$ in AQCD is then
${\cal F}(\sigma)_{\rm res.}=\int_{0}^{\infty}\frac{dt}{t}F_{{{\cal D}^{(1)}}}(t){\widetilde{\mathfrak{H}}}_{\nu_{0}}(te^{-{\widetilde{K}}_{e}}\sigma),$ (93)
where the timelike analog ${\widetilde{\mathfrak{H}}}_{\nu_{0}}(\sigma)$ of the generalised logarithmic derivative coupling ${\widetilde{\mathcal{A}}}_{\nu_{0}}(Q^{2})$ in AQCD is given in Eqs. (76).
Since the expansion of ${\widetilde{B}}[{{\cal D}^{(1)}}](u)$ in powers of $u$ generates the coefficients ${\widetilde{d}}_{n}(1;\kappa)$ (with $\kappa=1$), cf. Eq. (13), the knowledge of the four parameters ${\widetilde{K}}_{q}^{(1)}$ ($q=e,+,0,$-) needed to obtain the characteristic function $F_{{{\cal D}^{(1)}}}(t)$ is equivalent to the knowledge of the first four coefficients ${\widetilde{d}}_{n}(1;\kappa)$ ($n=0,1,2,3$; with $\kappa=1$). In order to obtain these four coefficients, the question is how they are related to the aforementioned (and known) coefficients $f_{n}(\nu_{0};\kappa)$ ($n=0,1,2,3$; with $\kappa=1$). These relations are obtained as follows.
The knowledge of the first four coefficients $f_{n}(\nu_{0};\kappa)$ ($\kappa=1$), Eqs. (81)-(82), gives us ${\widetilde{f}}_{n}(\nu_{0};\kappa)$ ($\kappa=1$) via the relations of the form Eq. (3) [^27] These relations are valid for the expansion coefficients of spacelike and timelike observables.
${\widetilde{f}}_{n}(\nu_{0},\kappa)=\sum_{s=0}^{n}{\widetilde{k}}_{n-s}(\nu_{0}+s)f_{s}(\nu_{0};\kappa),$ (94)
where we recall that the coefficients ${\widetilde{k}}_{n-s}(\nu_{0}+s)$ are given in Ref. GCAK . The coefficients ${\widetilde{d}}_{n}(\nu_{0};\kappa)$ (with $\kappa=1$; $0\leq n\leq 3$) are then obtained by inverting the first four equations (65). And finally, the coefficients ${\widetilde{d}}_{n}(1;\kappa)$ (with $\kappa=1$; $0\leq n\leq 3$) are then obtained by the rescaling relations (9). This then allows us to obtain the four coefficients ${\widetilde{K}}_{q}^{(1)}$ appearing in the modified Borel Eq. (91) by expanding this expression up to $u^{3}$, where we recall that ${\widetilde{B}}[{{\cal D}^{(1)}}](u)=$ $\sum_{n\geq 0}{\widetilde{d}}_{n}(1;1)u^{n}/(n!\beta_{0}^{n})$, and by equating the coefficients at $u^{n}$ ($n=0,1,2,3$).
If the considered observable (to be resummed) is spacelike, ${\cal D}(Q^{2})$, then the procedure is somewhat shorter. Namely, the knowledge of the first four coefficients $d_{n}(\nu_{0};\kappa)$ (with $\kappa=1$) gives us the first four coefficients ${\widetilde{d}}_{n}(\nu_{0};\kappa)$ (with $\kappa=1$) by using the relations (3), and the coefficients ${\widetilde{d}}_{n}(1;\kappa)$ (with $\kappa=1$; $0\leq n\leq 3$) are obtained by the rescaling relations (9). This allows us to obtain the four parameters ${\widetilde{K}}_{q}^{(1)}$ appearing in the modified Borel (91), and thus the characteristic function $F_{{{\cal D}^{(1)}}}(t)$ of Eq. (92) that replaces $F_{{{\cal D}^{(1)}}}(t)_{(p,{\widetilde{s}})}$ in the resummations Eqs. (79b) and Eqs. (78b) for ${\cal D}(Q^{2})$.
[…]
In Table 3 we present the values of the first four (exactly known) expansion coefficents $f_{n}(\nu_{0};\kappa)$, $d_{n}(\nu_{0};\kappa)$ and $d_{n}(1;\kappa)$ (all for $\kappa=1$), for $n_{f}=3,4$.
Table 3: The various coefficients appearing in our considered quantities, for $\kappa=1$ and for $n_{f}=3$ (and in parentheses for $n_{f}=4$). The coefficients ${\bar{f}}_{n}$ are in the $\overline{\rm MS}$ scheme [which has: ${\bar{c}}_{2}=4.47106(3.04764)$; ${\bar{c}}_{3}=20.9902(15.066)$; ${\bar{c}}_{4}=56.5876(27.3331)$]. All other coefficients are in the LMM scheme [which has: $c_{2}=9.29703(6.36801)$; $c_{3}=71.4538(50.8025)$; $c_{4}=201.843(74.2128)$]. All the values for $n\leq 3$ are exact. The values for $n\geq 4$ come from from the modified Borel Eq. (91). The values $f_{n}$ and ${\bar{f}}_{n}$ were not estimated beyond $n=4$ because we do not have the expressions for the coefficients $k_{n-s}(\nu_{0}+s)$ [appearing in the relation (4), valid also for $f_{n}$’s] for ns>4n-s>4.
$n$ ${\bar{f}}_{n}(\nu_{0})$ $f_{n}(\nu_{0})$ ${\widetilde{f}}_{n}(\nu_{0})$ ${\widetilde{d}}_{n}(\nu_{0})$ ${\widetilde{d}}_{n}(1)$
0 1 (1) 1 (1) 1 (1) 1 (1) 1 (1)
1 2.15741 (2.11227) 2.15741 (2.11227) 2.48619 (2.40388) 2.48619 (2.40388) 7.45857 (6.67744)
2 16.8999 (14.6762) 15.2913 (13.4809) 14.8747 (12.9312) 18.5758 (16.4267) 83.5909 (67.1027)
3 193.42 (150.119) 171.127 (134.148) 131.911 (104.206) 196.322 (159.291) 1135.86 (827.158)
4 3485.26 (2533.53) 3216.58 (2367.68) 2235.09 (1696.96) 3366.32 (2567.56) 23371.8 (15872.3)
5 - - 47196.3 (33189.2) 68486.8 (48121.1) 548647. (341144.)
6 - - $1.15082\times 10^{6}$ (746631.) $1.73121\times 10^{6}$ ($1.12989\times 10^{6}$) $1.56023\times 10^{7}$ ($8.96651\times 10^{6}$)
7 - - $3.38875\times 10^{7}$ ($2.04332\times 10^{7}$) $5.12524\times 10^{7}$ ($3.09886\times 10^{7}$) $5.10527\times 10^{8}$ ($2.70665\times 10^{8}$)
8 - - $1.14941\times 10^{9}$ ($6.42552\times 10^{8}$) $1.74758\times 10^{9}$ ($9.79388\times 10^{8}$) $1.89902\times 10^{10}$ ($9.29815\times 10^{9}$)
In Table 4 we present the parameters ${\widetilde{K}}_{q}$ of the modified Borel ${\widetilde{B}}[{{\cal D}^{(1)}}](u)$ Eq. (91) (at $\kappa=1$), in the mentioned LMM scheme, for $n_{f}=3,4$. We recall that these parameters then determine the characteristic function $F_{{{\cal D}^{(1)}}}(t)$ (92).
Table 4: The parameters ${\widetilde{K}}_{q}$ appearing in the modified Borel transform ${\widetilde{B}}[{{\cal D}^{(1)}}](u)$ Eq. (91), for $n_{f}=3$ and $n_{f}=4$ (in the LMM scheme, and for $\kappa=1$). The corresponding scheme parameter values $c_{j}$ ($j=2,3$) and the power index $\nu_{0}$ are included.
$n_{f}$ ${\widetilde{K}}_{e}$ $\widetilde{K}$$_{+}$ ${\widetilde{K}}_{0}$ $\widetilde{K}$$_{-}$ $c_{2}$ $c_{3}$ $\nu_{0}(n_{f})$
$n_{f}=3$ -2.23644 8.45107 -16.6796 8.39384 9.29703 71.4538 $1/3$
$n_{f}=4$ -2.36840 7.50692 -14.8007 7.45394 6.36801 50.8025 $9/25$
We then evaluate the timelike quantity of interest ${\cal F}(\sigma)$ with the resummation (93), for $n_{f}=3,4$. In particular, we evaluate it: (a) at $\sigma=m_{c}^{2}$ (with $n_{f}=3$) where $m_{c}=(1.67\pm 0.07)$ GeV is the pole mass of charm quark PDG2024 ; (b) at $\sigma=m_{b}^{2}$ (with $n_{f}=4$) where $m_{b}=(4.78\pm 0.06)$ GeV is the pole mass of the bottom quark PDG2024 .
Furthermore, we also evaluate these quantities with the naive perturbation approach (truncated perturbation series (TPS)) in powers of the $\overline{\rm MS}$ pQCD coupling:
${\cal F}(\sigma)^{{\rm TPS[N]};\overline{\rm MS}}={\bar{a}}(\sigma)^{\nu_{0}}+\sum_{j=1}^{N-1}{\bar{f}}_{j}{\bar{a}}(\sigma)^{\nu_{0}+j},$ (95)
where the bars indicate that these are the quantities in the $\overline{\rm MS}$ scheme, either with $n_{f}=3$ or $n_{f}=4$.
In Fig. 2 we present the results for the resummed ${\cal F}(\sigma)$, Eq. (78c), for $n_{f}=3$, as a function of $\sigma$ in the interval $0.1\ {\rm GeV}^{2}<\sigma<3\ {\rm GeV}^{2}$.
[…]
Figure 3: The resummed values of ${\cal F}(\sigma)$ as in Fig. 2, when (a) the IR-threshold scale $M_{1}$ is varied and $\alpha_{s}^{\overline{\rm MS}}(M_{Z}^{2})=0.1180$; (b) when $\alpha_{s}^{\overline{\rm MS}}(M_{z}^{2})$ value is varied and $M_{1}=0.150$ GeV.
The results depend significantly on the variation of the (representative) nonperturbative physics parameter $M_{1}$ [the IR-threshold of the spectral function of the coupling $\mathcal{A}$}, cf. Eq. (73)], especially at low $\sigma$ values. Nonetheless, as we will see, the pQCD results (at $\sigma=m_{c}^{2}$) possess a significantly larger uncertainty originating from the IR-renormalon ($u=1/2$) ambiguity.
[…]
${\cal F}(m_{c}^{2})_{\rm res.}$ $=$ $0.6365^{-0.0058}_{+0.0060}(m_{c})^{+0.0108}_{-0.0105}(\alpha_{s})^{+0.0243}_{-0.0245}(M_{1})$ (96a)
$=$ $0.6365\pm 0.0273.$ (96b)
In Eq. (96a) we denoted separately the uncertainties stemming from the variations of the pole mass $m_{c}=(1.67\pm 0.07)$ GeV, the variations of $\alpha_{s}^{\overline{\rm MS}}(M_{Z}^{2})=0.1180\pm 0.009$, and the variations of the (IR-)threshold scale $M_{1}=(0.150^{+0.100}_{-0.050})$ GeV. In Eq. (96b), the variations were added in quadrature.
[…]
${\cal F}(m_{b}^{2})_{\rm res.}$ $=$ $0.4792\mp 0.0010(m_{b})^{+0.0055}_{-0.0053}(\alpha_{s})^{+0.0061}_{-0.0065}(M_{1})$ (97a)
$=$ $0.4792^{+0.0084}_{-0.0083}.$ (97b)
The uncertainties at ’($m_{b}$)’ come from the uncertainty of the $b$ quark pole mass $m_{b}=(4.78\pm 0.06)$ GeV. We can see that the IR-uncertainties $(M_{1})$’ are relatively large in the case of ${\cal F}(m_{c}^{2})_{\rm res.}$ (with $n_{f}=3$), but less so for ${\cal F}(m_{b}^{2})_{\rm res.}$ (with $n_{f}=4$).
On the other hand, the naive pQCD TPS approach Eq. (95), in the $\overline{\rm MS}$ scheme, gives for $n_{f}=3$, at $\sigma=m_{c}^{2}$
${\cal F}(m_{c}^{2})^{\rm TPS[3]}$ $=$ $0.6604^{-0.0116}_{+0.0130}(m_{c})^{+0.0120}_{-0.0115}(\alpha_{s})\pm 0.0854({\rm TPS})$ (98a)
$=$ $0.6604^{+0.0872}_{-0.0869},$ (98b)
and for $n_{f}=4$, at $\sigma=m_{b}^{2}$
${\cal F}(m_{b}^{2})^{\rm TPS[4]}$ $=$ $0.4807\mp 0.0014(m_{b})\pm 0.0049(\alpha_{s})\pm 0.0184({\rm TPS})$ (99a)
$=$ $0.4807\pm 0.0191.$ (99b)
In both of these cases, the series was truncated at the smallest term (and including that term). This means that in the case of $\sigma=m_{c}^{2}$ the series was truncated at $N=3$ (the last term is $\sim{\bar{a}}^{\nu_{0}+2}$), and in the case of $\sigma=m_{b}^{2}$ it was truncated at $N=4$. The IR-uncertainties in these cases, ’(TPS)’, were taken to be these last (smallest) terms, which also indirectly reflect the IR-renormalon pQCD uncertainties ($u=1/2$) of these quantities. When we compare the ’(TPS)’ uncertainties in this pQCD approach for ${\cal F}(m_{c}^{2})$, Eq. (98a), with those of ’($M_{1}$)’ IR-uncertainties in Eq. (96a), we see that the latter are much smaller, and this despite choosing relatively large variations of the threshold scale $M_{1}$ around the pion mass value.
In principle, we could have evaluated ${\cal F}(\sigma)$ in the renormalon-motivated resummation with the (miniMOM) pQCD coupling, Eq. (79c). However, it turns out that this evaluation gives us significantly lower results; the central results are in this case: ${\cal F}(m_{c}^{2})_{\rm res.,pQCD}=0.4645$ and ${\cal F}(m_{b}^{2})_{\rm res.,pQCD}=0.4264$. The reason for this discrepancy with the above results lies probably in the significantly large Landau cut of the (miniMOM) pQCD coupling $a(Q^{2})$, namely this cut is $(0\leq)$$Q^{2}<1.27\ {\rm GeV}^{2}$ for $n_{f}=3$ and $(0\leq)$$Q^{2}<0.83\ {\rm GeV}^{2}$ for $n_{f}=4$. This then results in the same large Landau cuts for the coupling ${\widetilde{a}}_{\nu_{0}}(Q^{2})$, where [^29] The results for ${\cal F}(m_{q}^{2})_{\rm res.,pQCD}$ are up to four digits independent of whether or not we include the term $k_{4}(\nu_{0})a(Q^{2})^{\nu_{0}+4}$ in the sum Eq. (2a) that gives us ${\widetilde{a}}_{\nu_{0}}(Q^{2})$. we recall that the timelike pQCD coupling ${\widetilde{\mathfrak{h}}}_{\nu_{0}}(\sigma)$ appearing in the resummation integral (79c) is a contour integral of ${\widetilde{a}}_{\nu_{0}}(\sigma e^{i\phi})$ according to Eq. (55), and the value of this contour integral is for low $\sigma$ values certainly affected by the mentioned large Landau cuts.
We point out that the quantities ${\hat{C}}(m_{q})$ in the work of GandN , the scheme invariant factor of the Wilson coefficient of the chromomagnetic factor of the heavy-quark effective Lagrangian, are connected with our (“canonical”) function ${\cal F}(\sigma)$ in the following way:
${\hat{C}}(m_{q})=\pi^{\nu_{0}}{\cal F}(m_{q}^{2})\,$ (100)
where we choose to take for the number of (effectively) massless quarks $n_{f}=3$ when $m_{q}=m_{c}$, and $n_{f}=4$ when $m_{q}=m_{b}$. We note that $\nu_{0}$ depends on $n_{f}$, i.e., $\nu_{0}=\nu_{0}(n_{f})$ In particular, the ratios of these two functions at $m_{q}=m_{c}$ and $m_{q}=m_{b}$ is then
$\frac{{\hat{C}}(m_{b})}{{\hat{C}}(m_{c})}=\pi^{\nu_{0}(4)-\nu_{0}(3)}\frac{{\cal F}(m_{b}^{2})}{{\cal F}(m_{c}^{2})}.$ (101)
Here, $\nu_{0}(4)-\nu_{0}(3)=2/75$. The above results Eqs. (96)-(99) imply for the ratios the following values:
$\left(\frac{{\hat{C}}(m_{b})}{{\hat{C}}(m_{c})}\right)_{\rm res.}$ $=$ $0.7763^{+0.0071}_{-0.0073}(m_{c})\mp 0.0017(m_{b})^{-0.0196}_{+0.0204}(M_{1}\text{\&}\alpha_{s}),$ (102a)
$\left(\frac{{\hat{C}}(m_{b})}{{\hat{C}}(m_{c})}\right)_{\rm TPS}$ $=$ $0.7504^{+0.0134}_{-0.0145}(m_{c})^{-0.0021}_{+0.0022}(m_{b})^{-0.0608}_{+0.0787}({\rm TPS}\text{\&}\alpha_{s}).$ (102b)
In these ratios, we considered that the uncertainties coming from the IR regime $(M_{1})$’ or (TPS)] in the numerator and the denominator are completely correlated, as are the $(\alpha_{s})$’ uncertainties. These two types of uncertainties were then added in quadrature.
It turns out that this ratio is the leading order approximation for the following ratio of mass splitting between the ground-state pseudoscalar and vector mesons, in the bottom and charm quark systems GandN
$\frac{M_{B^{*}}^{2}-M_{B}^{2}}{M_{D^{*}}^{2}-M_{D}^{2}}=\frac{{\hat{C}}(m_{b})}{{\hat{C}}(m_{c})}\left[1+\Lambda_{\rm eff}\left(\frac{1}{m_{c}}-\frac{1}{m_{b}}\right)+\ldots\right],$ (103)
and $\Lambda_{\rm eff}$ in the subleading terms is a combination of the hadronic parameters. This ratio of mass splitting is $0.8776$, using the data PDG2024 . We now use in this relation the results (102) and we extract the value of this hadronic parameter
$\Lambda_{\rm eff}$ $=$ $\left(0.335^{-0.006}_{+0.005}(m_{c})\mp 0.004(m_{b})^{+0.075}_{-0.074}(M_{1}\text{\&}\alpha_{s})\right)\ {\rm GeV}$ (104a)
$=$ $\left(0.335\pm 0.075\right)\ {\rm GeV}\qquad{\rm(res)},$
$\Lambda_{\rm eff}$ $=$ $\left(0.435^{-0.027}_{+0.028}(m_{c})\pm 0.006(m_{b})^{+0.265}_{-0.285}({\rm TPS}\text{\&}\alpha_{s})\right)\ {\rm GeV}$ (104b)
$=$ $\left(0.435^{+0.266}_{-0.286}\right)\ {\rm GeV}\qquad{\rm(TPS)}.$
[…]
To improve the $\overline{\rm MS}$ TPS result in the future, it is desirable to have a parametric control on the renormalon contributions in the Hyperasymptotic approximation AP1 ; AP2 .
VI Conclusions
A renormalon-motivated resummation of QCD observables, very convenient in QCD formulations where the running coupling is free from the Landau singularities (AQCD: holomorphic QCD), was first developed renmod for the evaluation of (observable) quantities whose perturbation expansion has integer powers of the coupling.
In order to perform the evaluation, we have to know the renormalon structure of the considered quantity, and the first few coefficients of the perturbation expansion.
In this work we extend this formalism to the QCD observables, either spacelike or timelike, whose perturbation expansion has in general noninteger powers of the coupling.
As an example of specific application, we evaluated with this formalism the scheme-invariant factor ${\hat{C}}(m_{q})$ of the Wilson coefficient of the chromomagnetic operator in the heavy-quark effective Lagrangian, and related quantities. We used in our approach a running coupling ${\mathcal{A}}(Q^{2})$ free from the Landau singularities. The IR-behaviour of the spectral function $\rho_{{\mathcal{A}}}(\sigma)$ of the (holomorphic) coupling ${\mathcal{A}}(Q^{2})$ is parametrised with a sum of three Dirac delta functions, and we impose on ${\mathcal{A}}(Q^{2})$ at low $|Q^{2}|\lesssim 1\ {\rm GeV}^{2}$ the behaviour as suggested by large-volume lattice calculations. The IR-uncertainty was parametrised by varying significantly the (IR-)threshold scale $M_{1}^{2}$ of $\rho_{{\mathcal{A}}}(\sigma)$ around the value of the squared pion mass $m^{2}_{\pi}$. The obtained values of the factor functions ${\hat{C}}(m_{c})$ and ${\hat{C}}(m_{b})$ are somewhat different from those of the naive evaluation with the truncated perturbation series (TPS); however, the IR-uncertainties of the obtained results turn out to be much smaller in our approach than in the TPS approach, especially for the low-energy quantity ${\hat{C}}(m_{c})$.
We hope that this formalism can be useful in theoretical evaluations of several other low-energy QCD observables.
Acknowledgements.
This work was supported in part by FONDECYT (Chile) Grants No. 1220095 (G.C.) and 1240329 (C.A.).
Appendix A Proof of Theorem 2
For $d(\nu_{0};\kappa)$ we have, using (25), the expansion (29) and the asymptotic formula (30)
$d_{n}(\nu_{0};\kappa)$ $=$ $\frac{{\cal K}(\kappa)p^{-s_{0}}}{\Gamma(s_{0})}\Gamma(s_{0}+n)\left(\frac{\beta_{0}}{p}\right)^{n}\left(1+{\cal O}(1/n)\right)$ (105a)
$=$ $\frac{{\cal K}(\kappa)p^{-s_{0}}}{\Gamma(s_{0})}n!n^{s_{0}-1}\left(\frac{\beta_{0}}{p}\right)^{n}\left(1+{\cal O}(1/n)\right).$ (105b)
This is a situation similar to the one considered in Appendix B of Ref. GCAK . There, $s_{0}=\nu_{0}$ (while here $s_{0}\not=\nu_{0}$ in general), $\beta_{0}/p\mapsto b$ and $d_{n}(\nu_{0};\kappa)\mapsto{\cal F}_{n}$. Further, there we have no ${\cal O}(1/n)$ corrections; nonetheless, since the ${\cal O}(1/n)$-terms will be kept here undetermined throughout, we can follow the proof in Appendix B of GCAK , with the only modifications arising now from the fact that $s_{0}\not=\nu_{0}$.
The idea of the proof of Theorem 2 is to find the ratio between the coefficients ${\widetilde{d}}_{n}(\nu_{0};\kappa)$ and $d_{n}(\nu_{0},\kappa)$ for large $n$. For this, the relations (3) at large $n$ will be investigated here.303030 The notational conventions for the $\beta$-function coefficients $\beta_{j}$ and $c_{j}$ are given in Eqs. (6) where, as always, $a=\alpha_{s}/\pi$.
On the right-hand side (RHS) of Eq. (3) we have, for $s=n$, ${\widetilde{k}}_{0}(\nu_{0}+s)=1$, and thus
$\frac{{\widetilde{k}}_{0}(\nu_{0}+n)d_{n}(\nu_{0};\kappa)}{d_{n}(\nu_{0};\kappa)}=1.$ (106)
From now on, for convenience, we will use the notations
$b\equiv\frac{\beta_{0}}{p};\quad K\equiv{\cal K}(\kappa)p^{-s_{0}}.$ (107)
With this notation, the behaviour of $d_{n}(\nu_{0};\kappa)$ in the considered case, Eq. (105a), can be written as
$d_{n}(\nu_{0};\kappa)=K\frac{\Gamma(s_{0}+n)}{\Gamma(s_{0})}b^{n}\left(1+{\cal O}(1/n)\right).$ (108)
Further, we will repeatedly use Eqs. (A.17)-(A.21) of Ref. GCAK for the coefficients ${\widetilde{k}}_{m}(\nu)$, and Eq. (A.13) for the function ${\widetilde{Z}}_{m}(\nu)$ appearing in those expressions
${\widetilde{Z}}_{m}(\nu)=\Gamma(\nu+1)\left(\frac{d}{dx}\right)^{m}\left(\frac{\Gamma(1-x)}{\Gamma(\nu+1-x)}\right){\Bigg{|}}_{x=0}.$ (109)
Then, on the RHS of Eq. (3) we have for s=n1s=n-1
${\widetilde{k}}_{1}(\nu_{0}+n-1)d_{n-1}(\nu_{0};\kappa)$ $=$ $-c_{1}\nu\left[{\widetilde{Z}}_{1}(\nu)-1\right]\frac{\Gamma(s_{0}+n-1)}{\Gamma(s_{0})}b^{n-1}\left(1+{\cal O}(1/n)\right)K{\Big{|}}_{\nu=\nu_{0}+n-1}\;\Rightarrow$ (110a)
$\frac{{\widetilde{k}}_{1}(\nu_{0}+n-1)d_{n-1}(\nu_{0};\kappa)}{d_{n}(\nu_{0};\kappa)}$ $=$ $-c_{1}\frac{1}{b}\left(\frac{d}{dx}-1\right)\left(\frac{\Gamma(1-x)\Gamma(\nu_{0}+n)}{\Gamma(\nu_{0}+n-x)}\right){\Bigg{|}}_{x=0}\times\left(1+{\cal O}(1/n)\right).$ (110b)
The second identity (110b) is obtained by the use of the identity (109) for ${\widetilde{Z}}_{1}$ and the known property of the Gamma function, $\Gamma(z+1)=z\Gamma(z)$.
Similarly, on the RHS of Eq. (3) we have for s=n2s=n-2
$\frac{{\widetilde{k}}_{2}(\nu_{0}+n-2)d_{n-2}(\nu_{0};\kappa)}{d_{n}(\nu_{0};\kappa)}$ $=$ $\nu(\nu+1)\left[-c_{2}\frac{(\nu-1)}{2(\nu+1)}+\frac{1}{2}c_{1}^{2}\left({\widetilde{Z}}_{2}(\nu)-2{\widetilde{Z}}_{1}(\nu+1)+1\right)\right]$ (111a)
$\times\frac{\Gamma(s_{0}+n-2)}{\Gamma(s_{0}+n)}\frac{1}{b^{2}}\left(1+{\cal O}(1/n)\right)$
$=$ $\frac{1}{2b^{2}}\left[c_{1}^{2}\left(\frac{d}{dx}-1\right)^{2}-c_{2}\right]\left(\frac{\Gamma(1-x)\Gamma(\nu_{0}+n)}{\Gamma(\nu_{0}+n-x)}\right){\Bigg{|}}_{x=0}\times\left(1+{\cal O}(1/n)\right).$ (111b)
In arriving to Eq. (111b), we in addition used the relation
${\widetilde{Z}}_{2}(\nu_{0}+n-2)={\widetilde{Z}}_{2}(\nu_{0}+n-1)\left(1+{\cal O}(1/n)\right).$ (112)
Analogously, on the RHS of Eq. (3) we obtain for s=n3s=n-3
$\frac{{\widetilde{k}}_{3}(\nu_{0}+n-3)d_{n-3}(\nu_{0};\kappa)}{d_{n}(\nu_{0};\kappa)}$ $=$ $\frac{1}{6b^{3}}\left[-c_{1}^{3}\left(\frac{d}{dx}-1\right)^{3}+3c_{1}c_{2}\left(\frac{d}{dx}-\frac{1}{6}\right)-\frac{1}{2}c_{3}\right]\left(\frac{\Gamma(1-x)\Gamma(\nu_{0}+n)}{\Gamma(\nu_{0}+n-x)}\right){\Bigg{|}}_{x=0}$ (113)
$\times\left(1+{\cal O}(1/n)\right).$
and for s=n4s=n-4
$\frac{{\widetilde{k}}_{4}(\nu_{0}+n-4)d_{n-4}(\nu_{0};\kappa)}{d_{n}(\nu_{0};\kappa)}$ $=$ $\frac{1}{24b^{4}}{\Bigg{\{}}c_{1}^{4}\left(\frac{d}{dx}-1\right)^{4}-6c_{1}^{2}c_{2}\left[\left(\frac{d}{dx}-\frac{1}{6}\right)^{2}+\frac{31}{36}\right]+2c_{1}c_{3}\left(\frac{d}{dx}+\frac{1}{6}\right)$ (114)
$+\frac{(13c_{2}^{2}-c_{4})}{3}{\Bigg{\}}}\left(\frac{\Gamma(1-x)\Gamma(\nu_{0}+n)}{\Gamma(\nu_{0}+n-x)}\right){\Bigg{|}}_{x=0}\times\left(1+{\cal O}(1/n)\right).$
Appendix A.1
The rest of this proof consists of two parts. The first part of the rest of this proof considers the effects of the terms $c_{1}$ ($\beta_{1}/\beta_{0}$), regarding $c_{n}=0$ ($n=2,3,\ldots$).
In this part we now follow entirely Appendix B.1 of Ref. GCAK . This is so because the asymptotic results of the previous part of the present Appendix, Eqs. (110b), …, (114), are the same as in App. B of GCAK (although we have here $s_{0}\not=\nu_{0}$ in general).
When taking the terms in powers of $c_{1}$ appearing in Eqs. (110b), …, (114), we obtain
${\widetilde{d}}_{n}(\nu_{0};\kappa)^{(c_{1})}$ $=$ $d_{n}(\nu_{0};\kappa)\sum_{m=0}^{n}\frac{(-1)^{m}c_{1}^{m}}{b^{m}m!}\left(\frac{d}{dx}-v\right)^{m}\left(\frac{\Gamma(1-x)\Gamma(\nu_{0}+n)}{\Gamma(\nu_{0}+n-x)}\right){\Bigg{|}}_{x=0}\times\left(1+{\cal O}(1/n)\right),$ (115)
where $v=1$ and the superscript $(c_{1})$’ denotes that only those terms are taken into account that are nonzero when $c_{j}=0$ for $j\geq 2$. The last term in Eq. (115) we represent by using the integral form of the mathematical beta-function $B(u,w)$
$\frac{\Gamma(1-x)\Gamma(\nu_{0}+n)}{\Gamma(\nu_{0}+n-x)}$ $=$ $(\nu_{0}+n-1)B(1-x,\nu_{0}+n-1)=(\nu_{0}+n-1)\int_{0}^{1}dyy^{-x}(1-y)^{\nu_{0}+n-2}.$ (116)
Application of $(d/dx-v)^{m}$ to the identity (116) means that this operator is applied to the factor $y^{-x}=\exp(-x\ln y)$, and this leads, after some algebra, to the identity
$\left(\frac{d}{dx}-v\right)^{m}\left(\frac{\Gamma(1-x)\Gamma(\nu_{0}+n)}{\Gamma(\nu_{0}+n-x)}\right){\Bigg{|}}_{x=0}$ $=$ $(\nu_{0}+n-1)\int_{0}^{1}dy(-\ln y-v)^{m}(1-y)^{\nu_{0}+n-2}.$ (117)
When we use this identity in Eq. (115), we obtain
${\widetilde{d}}_{n}(\nu_{0};\kappa)^{(c_{1})}$ $=$ $d_{n}(\nu_{0};\kappa)(\nu_{0}+n-1)\int_{0}^{1}dy(1-y)^{\nu_{0}+n-2}\sum_{m=0}^{n}\frac{(\ln y+v)^{m}c_{1}^{m}}{b^{m}m!}{\Big{|}}_{v=1}\left(1+{\cal O}(1/n)\right)$ (118a)
$\approx$ $d_{n}(\nu_{0};\kappa)(\nu_{0}+n-1)\exp(vc_{1}/b)|_{v=1}\int_{0}^{1}dyy^{c_{1}/b}(1-y)^{\nu_{0}+n-2}\left(1+{\cal O}(1/n)\right)$ (118b)
$=$ $d_{n}(\nu_{0};\kappa)\exp(c_{1}/b)\frac{\Gamma\left(1+\frac{c_{1}}{b}\right)\Gamma(\nu_{0}+n)}{\Gamma\left(\nu_{0}+n+\frac{c_{1}}{b}\right)}\left(1+{\cal O}(1/n)\right).$ (118c)
In the step from Eq. (118a) to (118b) we replaced the sum $\sum_{m=0}^{n}$ by $\sum_{m=0}^{\infty}$, and in Eq. (118c) we explicitly used $v=1$. It can be checked numerically that this approximation results in relative errors which diminish with increasing $n$ significantly faster than ${\cal O}(1/n)$.
If we use on the RHS of Eq. (118c) the asymptotic formula (30) for the Gamma functions, we immediately obtain
${\widetilde{d}}_{n}(\nu_{0};\kappa)^{(c_{1})}$ $=$ $d_{n}(\nu_{0};\kappa)e^{c_{1}/b}\Gamma\left(1+\frac{c_{1}}{b}\right)n^{-c_{1}/b}\left(1+{\cal O}(1/n)\right)$ (119)
$=$ $d_{n}(\nu_{0};\kappa)e^{p\beta_{1}/\beta_{0}^{2}}\Gamma\left(1+\frac{p\beta_{1}}{\beta_{0}^{2}}\right)n^{-p\beta_{1}/\beta_{0}^{2}}\left(1+{\cal O}(1/n)\right),$
where we used in the last identity the explicit expression for $b$ Eq. (107) and $c_{1}=\beta_{1}/\beta_{0}$. The obtained asymptotic relation (119), when combined with the asymptotic expression (105b), then gives
${\widetilde{d}}_{n}(\nu_{0};\kappa)^{(c_{1})}$ $=$ $\frac{{\cal K}(\kappa)p^{-s_{0}}e^{p\beta_{1}/\beta_{0}^{2}}\Gamma(1+p\beta_{1}/\beta_{0}^{2})}{\Gamma(s_{0})}n!n^{(s_{0}-p\beta_{1}/\beta_{0}^{2})-1}\left(\frac{\beta_{0}}{p}\right)^{n}\left(1+{\cal O}(1/n)\right).$ (120)
As shown above around Eq. (105b), the asymptotic expression (105b) for $d_{n}(\nu_{0};\kappa)$ is equivalent to the form Eq. (25) of the Borel $B({\cal D}](u;\kappa)$. Comparing the obtained asymptotic expression (120) for ${\widetilde{d}}_{n}(\nu_{0};\kappa)^{(c_{1})}$ with that of $d_{n}(\nu_{0};\kappa)$ Eq. (105b), we see that the modified Borel ${\widetilde{B}}^{(\nu_{0})}[{\cal D}](u;\kappa)$ has the same structure, with the difference that the index $s_{0}$ is replaced by ${\widetilde{s}}_{0}=s_{0}-p\beta_{1}/\beta_{0}^{2}$.
This proves Theorem 2 for the case when the effects of the higher beta-coeffiicents are neglected (i.e., $c_{2}=c_{3}=\cdots=0$). [^31] The radiative corrections ${\cal O}(1/n)$ in the result (120) clearly correspond to $s_{0}\mapsto(s_{0}-1)$, i.e., in the modified Borel ${\widetilde{B}}^{(\nu_{0})}[{\cal D}](u;\kappa)$ they correspond to the relative corrections ${\cal O}(p-u)$, cf. Eq. (26). These corrections are in general nonzero even if the relative corrections ${\cal O}(p-u)$ in the Borel $B^{(\nu_{0})}[{\cal D}](u;\kappa)$ are taken to be zero.
Appendix A.2
Now we extend the proof of Theorem 2 to the case when $c_{2},c_{3},\ldots\not=0$. For this, we follow the steps formulated in Appendix B.2 of Ref. GCAK . Some of the steps involve “educated guess” extrapolations (especially for the cases $c_{3}\not=0$ and $c_{4}\not=0$).
First the terms that are nonzero when $c_{2}\not=0$, in Eqs. (106)-(114)
$\frac{{\widetilde{d}}_{n}(\nu_{0};\kappa)^{(c_{2})}}{d_{n}(\nu_{0};\kappa)}$ $=$ $\left\{-\frac{c_{2}}{2b^{2}}+\frac{c_{1}c_{2}}{2b^{3}}\left(\frac{d}{dx}-\frac{1}{6}\right)-\frac{c_{1}^{2}c_{2}}{4b^{4}}\left(\frac{d}{dx}-\frac{1}{6}\right)^{2}+\ldots\right\}\left(\frac{\Gamma(1-x)\Gamma(\nu_{0}+n)}{\Gamma(\nu_{0}+n-x)}\right){\Bigg{|}}_{x=0}.$ (121)
$\times\left(1+{\cal O}(1/n)\right).$
Here we excluded the curly brackets the term $(31/36)(c_{1}^{2}c_{2})/(4b^{4})$, we will combine it with the term $\sim c_{4}/b^{4}$ later. The pattern that we see in Eq. (121) gives
$\frac{{\widetilde{d}}_{n}(\nu_{0};\kappa)^{(c_{2})}}{d_{n}(\nu_{0};\kappa)}$ $=$ $-\frac{c_{2}}{2b^{2}}\sum_{m=0}^{n-2}\frac{(-c_{1})^{m}}{m!b^{m}}\left(\frac{d}{dx}-\frac{1}{6}\right)^{m}\left(\frac{\Gamma(1-x)\Gamma(\nu_{0}+n)}{\Gamma(\nu_{0}+n-x)}\right){\Bigg{|}}_{x=0}\left(1+{\cal O}(1/n)\right).$ (122)
[…]
$\frac{{\widetilde{d}}_{n}(\nu_{0};\kappa)^{(c_{2})}}{d_{n}(\nu_{0};\kappa)}$ $=$ $-\frac{c_{2}}{2b^{2}}e^{(1/6)c_{1}/b}\frac{\Gamma(1+c_{1}/b)\Gamma(\nu_{0}+n)}{\Gamma(\nu_{0}+n+c_{1}/b)}\left(1+{\cal O}(1/n)\right)$
$=$ $-\frac{c_{2}}{2b^{2}}e^{(1/6)c_{1}/b}\Gamma(1+c_{1}/b)n^{-c_{1}/b}\left(1+{\cal O}(1/n)\right)\;\Rightarrow$
${\widetilde{d}}_{n}(\nu_{0};\kappa)^{(c_{2})}$ $=$ $-\frac{c_{2}}{2b^{2}}{\widetilde{d}}_{n}(\nu_{0};\kappa)^{(c_{1})}e^{-(5/6)(c_{1}/b)}\left(1+{\cal O}(1/n)\right),$ (123)
where in the last identity we used the result (119). This identity is important, because it means that the effects of $c_{2}\not=0$ only rescale the full ${\widetilde{d}}_{n}(\nu_{0};\kappa)$ [in comparison to ${\widetilde{d}}_{n}(\nu_{0};\kappa)^{(c_{1})}$] by a factor that is independent of $n$ (in the leading order of large $n$). Below we will see that the effects of $c_{3}\not=0$ and $c_{4}\not=0$ result in a similar kind of rescaling of ${\widetilde{d}}_{n}(\nu_{0};\kappa)$.
The terms in Eqs. (106)-(114) that are nonzero when $c_{3}\not=0$, are
$\frac{{\widetilde{d}}_{n}(\nu_{0};\kappa)^{(c_{3})}}{d_{n}(\nu_{0};\kappa)}$ $=$ $\left\{-\frac{c_{3}}{12b^{3}}+\frac{c_{1}c_{3}}{12b^{4}}\left(\frac{d}{dx}+\frac{1}{6}\right)+\ldots\right\}\left(\frac{\Gamma(1-x)\Gamma(\nu_{0}+n)}{\Gamma(\nu_{0}+n-x)}\right){\Bigg{|}}_{x=0}\left(1+{\cal O}(1/n)\right)$ (124a)
$=$ $-\frac{c_{3}}{12b^{3}}\sum_{m=0}^{n-3}\frac{(-c_{1})^{m}}{m!b^{m}}\left(\frac{d}{dx}+\frac{1}{6}\right)^{m}\left(\frac{\Gamma(1-x)\Gamma(\nu_{0}+n)}{\Gamma(\nu_{0}+n-x)}\right){\Bigg{|}}_{x=0}\left(1+{\cal O}(1/n)\right).$ (124b)
Following the same steps as before, we obtain
${\widetilde{d}}_{n}(\nu_{0};\kappa)^{(c_{3})}$ $=$ $-\frac{c_{3}}{12b^{3}}{\widetilde{d}}_{n}(\nu_{0};\kappa)^{(c_{1})}e^{-(7/6)(c_{1}/b)}\left(1+{\cal O}(1/n)\right).$ (125)
When we consider the remaining terms, $\sim 1/b^{4}$, in Eqs. (106)-(114), we obtain
$\frac{{\widetilde{d}}_{n}(\nu_{0};\kappa)^{(c_{4})}}{d_{n}(\nu_{0};\kappa)}$ $=$ $\frac{(-1)}{144b^{4}}(2c_{4}-26c_{2}^{2}+31c_{1}^{2}c_{2})\times 1\times\left(1+{\cal O}(1/n)\right).$ (126)
When we add additional terms $\sim 1/b^{5}$, $\sim 1/b^{6}$, etc., we expect to obtain analogously an expression of the form
$\frac{{\widetilde{d}}_{n}(\nu_{0};\kappa)^{(c_{4})}}{d_{n}(\nu_{0};\kappa)}$ $=$ $\frac{(-1)}{144b^{4}}(2c_{4}-26c_{2}^{2}+31c_{1}^{2}c_{2})e^{-\beta^{(4)}c_{1}/b}\frac{\Gamma(1+c_{1}/b)\Gamma(\nu_{0}+n)}{\Gamma(\nu_{0}+n+c_{1}/b)}\left(1+{\cal O}(1/n)\right)$ (127a)
$=$ $\frac{(-1)}{144b^{4}}(2c_{4}-26c_{2}^{2}+31c_{1}^{2}c_{2})e^{-\beta^{(4)}c_{1}/b}\Gamma(1+c_{1}/b)n^{-c_{1}/b}\left(1+{\cal O}(1/n)\right),$ (127b)
where $\beta^{(4)}\sim 1$ (and probably $\beta^{(4)}>0$). Using in Eq. (127b) the results (119), this gives us
${\widetilde{d}}_{n}(\nu_{0};\kappa)^{(c_{4})}$ $=$ ${\widetilde{d}}_{n}(\nu_{0};\kappa)^{(c_{1})}\frac{(-1)}{144b^{4}}(2c_{4}-26c_{2}^{2}+31c_{1}^{2}c_{2})e^{-(\beta^{(4)}+1)c_{1}/b}\left(1+{\cal O}(1/n)\right).$ (128)
Combining the relations (123), (125) and (128), we obtain
${\widetilde{d}}_{n}(\nu_{0};\kappa)=\sum_{q=1}^{\infty}{\widetilde{d}}_{n}(\nu_{0};\kappa)^{(c_{q})}$ (129)
$=$ ${\widetilde{d}}_{n}(\nu_{0};\kappa)^{(c_{1})}\left[1-\frac{c_{2}}{2b^{2}}e^{-(5/6)(c_{1}/b)}-\frac{c_{3}}{12b^{3}}e^{-(7/6)(c_{1}/b)}-\frac{(2c_{4}-26c_{2}^{2}+31c_{1}^{2}c_{2})}{144b^{4}}e^{-(\beta^{(4)}+1)(c_{1}/b)}-\ldots\right],$
where we recall that $b=\beta_{0}/p$ and $c_{1}=\beta_{1}/\beta_{0}$. To obtain the excact value of the number $\beta^{(4)}$ in the exponent in Eq. (129), we would need to extend the explicit analysis of Eqs. (106)-(114) to the case of s=n5s=n-5, i.e., the inclusion of the explicit expression for ${\widetilde{k}}_{5}(\nu_{0}+n-5)$.
In the specific case considered in the main text of this work, we have $n_{f}=3$ and $p=1/2$, therefore $1/b=1/4.5=0.2222$ is small and the sum in Eq. (129) is expected to converge well.323232 For $p=1/2$ and $n_{f}=3$, the sum in the brackets on the RHS of Eq. (129) is: 10.07941-0.0794 0.0121-0.0121 $0.0005\times 0.7^{1+\beta^{(4)}}$.
What the result (129) shows is that the asymptotic behaviour of the coefficients ${\widetilde{d}}_{n}(\nu_{0};\kappa)$ is determined by the $c_{1}$-terms only, i.e., by the behaviour of ${\widetilde{d}}_{n}(\nu_{0};\kappa)^{(c_{1})}$ Eq. (120)
${\widetilde{d}}_{n}(\nu_{0};\kappa)$ $=$ ${\cal C}(\kappa)n!n^{(s_{0}-p\beta_{1}/\beta_{0}^{2})-1}\left(\frac{\beta_{0}}{p}\right)^{n}\left(1+{\cal O}(1/n)\right),$ (130)
where ${\cal C}(\kappa)$ is a constant independent of $n$. We recall that the asymptotic behaviour of $d_{n}(\nu_{0};\kappa)$, Eq. (105b), reflects the (i.e., is equivalent to the) form Eq. (25) of $B^{(\nu_{0})}[{\cal D}](u;\kappa)$ of Theorem 2. Our result Eq. (130) shows that ${\widetilde{d}}_{n}(\nu_{0};\kappa)$ has the same asymptotic behaviour, with the only difference that $s_{0}\mapsto s_{0}-p\beta_{1}/\beta_{0}^{2}$, which means that the corresponding modified Borel ${\widetilde{B}}^{(\nu_{0})}[{\cal D}](u;\kappa)$ of Eq. (26) has the power index ${\widetilde{s}}_{0}$ as claimed in Eq. (27) of the Theorem.
This concludes the proof of Theorem 2.
Appendix B Proof of Theorem 4
When the form of ${\widetilde{B}}[{{\cal D}^{(1)}}](u)$ is the renormalon term as given in Eq. (34), then the inverse Mellin transformation thereof, Eq. (16), gives us the corresponding characteristic function
$F_{{{\cal D}^{(1)}}}(t)$ $=$ $\frac{1}{2i}\int_{u=-i\infty}^{+i\infty}\frac{due^{u\ln t}}{(p-u)^{{\widetilde{s}}}}$, (131)
where we took $u_{0}=0$. We recall that $0<{\widetilde{s}}<1$. We introduce the change of the integration variable $z=iu$, and this gives
$F_{{{\cal D}^{(1)}}}(t)$ $=$ $\frac{1}{2}\int_{z=-\infty}^{+\infty}\frac{dze^{-iz\ln t}}{(iz+p)^{{\widetilde{s}}}}$, (132)
The cut of the integrand, in the complex $z$-plane, is for $(iz+p)\leq 0$, i.e., for $z=i|z|$ with $|z|\geq p$ (i.e., on the semiaxis along the positive imaginary axis).
When $t>1$, we have $\ln t=|\ln t|$, we can close the contour of integration of the integral (132) with the (large) semicircle in the lower half plane because of the exponential suppression there; Jordan Lemma ensures that the contribution along the semicircle of radius $R$, when $R\to\infty$, is zero. No singularities are enclosed, and this then gives us by the Cauchy theorem
$F_{{{\cal D}^{(1)}}}(t)=0\qquad(t>1).$ (133)
When $0\leq t<1$, we have $\ln t=-|\ln t|$, hence the integrand is exponentially suppressed if we close the contour with the (large) semicircle in the upper half plane. Jordan Lemma ensures that the contribution along the semicircle of radius $R$, when $R\to\infty$, is zero. Nonetheless, due to the aforementioned cut of the integrand along the positive imaginary axis ($z=i|z|$, $|z|\geq p$), we have to avoid this cut in order to apply the Cauchy theorem. The necessary contour is presented in Fig. 4.
Figure 4: The closed contour in the complex $z$-plane for the integral (132) for the case $0<t<1$. The limits $R\to+\infty$ and $\varepsilon\to+0$ are taken.
Since the closed contour avoids enclosing singularities, by Cauchy theorem the integral along the contour is zero
$\left\{\int_{-R}^{R}+\int_{{\cal C}_{+}(R)}+\int_{{\cal C}_{c}^{+}(\varepsilon)}+\int_{{\cal C}_{c}^{-}(\varepsilon)}+\int_{{\cal C}_{-}(\varepsilon)}\right\}\frac{dze^{+iz|\ln t|}}{2(iz+p)^{{\widetilde{s}}}}$. (134)
As mentioned, the contribution along (both parts of) the upper semicircle ${\cal C}_{+}(R)$ gives zero (when $R\to+\infty$) due to Jordan Lemma. We now show that the contribution around the small semicicle ${\cal C}_{-}(\varepsilon)$ or radius $\varepsilon$ (around the point $z=ip$) also gives zero (when $\varepsilon\to+0$). On that semicricle, we have $z=ip+\varepsilon\exp(i\phi)$ ($\pi<\phi<0$) and thus the contribution there is
$\frac{1}{2}\int_{{\cal C}_{-}(\varepsilon)}\frac{dze^{iz|\ln t|}}{(iz+p)^{{\widetilde{s}}}}$ $=$ $\frac{1}{2}i\varepsilon\int_{\phi=0}^{-\pi}\frac{d\phi e^{i\phi}\exp\left[-p|\ln t|+{\cal O}(\varepsilon)\right]}{(i\varepsilon e^{i\phi})^{{\widetilde{s}}}}$ (135)
$=$ $\varepsilon^{1-{\widetilde{s}}}\frac{(-1)}{2(1-{\widetilde{s}})}e^{-i{\widetilde{s}}\pi/2}e^{-p|\ln t|}\frac{1}{{\widetilde{s}}}\left(1+e^{i{\widetilde{s}}\pi}\right)\left(1+{\cal O}(\varepsilon)\right)\sim\varepsilon^{1-{\widetilde{s}}}$.
This goes clearly to zero when $\varepsilon\to+0$, because $0<{\widetilde{s}}<1$.
The contributions of the integral along both sides of the cut, ${\cal C}_{c}^{\pm}(\varepsilon)$ (cf. Fig. 4) can now be evaluated by introducing a new integration variable $w=iz+p$ and its absolute value $|w|$. The cut is now in the complex $w$-plane along the negative semiaxis, cf. Fig. 5.
Figure 5: The paths ${\cal C}_{c}^{\pm}(\varepsilon)$ in the complex $w$-plane, where $w=p+iz$. The limit $\varepsilon\to=)$ is taken.
We have dz=idwdz=-idw, and since on the two paths ${\cal C}_{c}^{\pm}(\varepsilon)$ we have $w=-|w|\pm i\varepsilon$ along this cut, therefore $dz=+id|w|$. Along ${\cal C}_{c}^{+}(\varepsilon)$ we have: $w=(-R+p)+i\varepsilon\to+i\varepsilon$, i.e., $w=|w|e^{+i(\pi-\varepsilon^{\prime})}$. Along ${\cal C}_{c}^{-}(\varepsilon)$ we have: $w=-i\varepsilon\to(-R+p)-i\varepsilon$, i.e., $w=|w|e^{-i(\pi-\varepsilon^{\prime})}$. Here, $\varepsilon\to+0$ and $\varepsilon^{\prime}\to+0$. The contributions are then
$\left\{\int_{{\cal C}_{c}^{+}(\varepsilon)}+\int_{{\cal C}_{c}^{-}(\varepsilon)}\right\}\frac{dze^{iz|\ln t|}}{2(iz+p)^{{\widetilde{s}}}}$ (136a)
$=$ $\frac{i}{2}{\Big{\{}}\int_{|w|=R-p}^{0}d|w|\frac{\exp\left(|w|e^{+i\pi}|\ln t|\right)e^{-p|\ln t|}}{\left(|w|e^{+i\pi}\right)^{{\widetilde{s}}}}+\int_{|w|=0}^{R-p}d|w|\frac{\exp\left(|w|e^{-i\pi}|\ln t|\right)e^{-p|\ln t|}}{\left(|w|e^{-i\pi}\right)^{{\widetilde{s}}}}{\Big{\}}}$
$=$ $\frac{i}{2}e^{-p|\ln t|}\int_{|w|=0}^{+\infty}d|w|\frac{e^{-|w||\ln t|}}{|w|^{{\widetilde{s}}}}\left(e^{i{\widetilde{s}}\pi}-e^{-i{\widetilde{s}}\pi}\right)=(-1)e^{-p|\ln t|}\sin({\widetilde{s}}\pi)\int_{|w|=0}^{+\infty}d|w||w|^{-{\widetilde{s}}}e^{-i|\ln t||w|}$ (136b)
$=$ $\frac{(-1)e^{-p|\ln t|}}{|\ln t|^{1-{\widetilde{s}}}}\sin({\widetilde{s}}\pi)\Gamma(1-{\widetilde{s}})=(-1)\frac{\pi e^{-p|\ln t|}}{\Gamma({\widetilde{s}})|\ln t|^{1-{\widetilde{s}}}}$. (136c)
In Eq. (136b) we took into account $R\to+\infty$. We now take into account this result in the Cauchy identity (134), and the fact that the contributions along the large semicircle ${\cal C}_{+}(R)$ and along the small semicircle ${\cal C}_{-}(\varepsilon)$ are zero [Eq. (135)]. This gives us the final result for $0<t<1$
$F_{{{\cal D}^{(1)}}}(t)_{(p,{\widetilde{s}})}$ $=$ $\frac{1}{2}\int_{-\infty}^{+\infty}\frac{dze^{+iz|\ln t|}}{(iz+p)^{{\widetilde{s}}}}=(-1)\left\{\int_{{\cal C}_{c}^{+}(\varepsilon)}+\int_{{\cal C}_{c}^{-}(\varepsilon)}\right\}\frac{dze^{+iz|\ln t|}}{2(iz+p)^{{\widetilde{s}}}}$ (137)
$=$ $\frac{\pi e^{-p|\ln t|}}{\Gamma({\widetilde{s}})|\ln t|^{1-{\widetilde{s}}}}=\frac{\pi t^{p}}{\Gamma({\widetilde{s}})(-\ln t)^{1-{\widetilde{s}}}}\quad(0\leq t\leq 1).$
We note that in this case $e^{-p|\ln t|}=e^{p\ln t}=t^{p}$. The result Eqs. (133) and (137) then give us the final result Eq. (35) for the characteristic function $F_{{{\cal D}^{(1)}}}(t)_{(p,{\widetilde{s}})}$.
This concludes the proof of Theorem 4.
The Theorem was proven for the case when the power index is $0<{\widetilde{s}}<1$. In Ref. renmod , the characteristic function was evaluated for the case of the IR renormalons with ${\widetilde{s}}=1$ and ${\widetilde{s}}=2$ (simple and double poles). It is interesting that the formula of Theorem 4 can be applied even in these cases ${\widetilde{s}}=1,2$.
Further, in the limit ${\widetilde{s}}\to 0$ we can take into account that
$\lim\limits_{{\widetilde{s}}\to 0}\pi\frac{(-1)}{{\widetilde{s}}}\left(\frac{1}{(p-u)^{{\widetilde{s}}}}-1\right)-\pi\ln p=\pi\ln\left(1-\frac{u}{p}\right).$ (138)
This means that in such a case we have to take in the result (137) the limit ${\widetilde{s}}\to 0$. Furthermore, $1/({\widetilde{s}}\Gamma({\widetilde{s}}))\to 1$ when ${\widetilde{s}}\to 0$. However, since such ${\widetilde{B}}[{{\cal D}^{(1)}}](u)_{(p,0)}$ does not produce the constant term $\sim u^{0}$ in its expansion, the corresponding ${{\cal D}^{(1)}}(Q^{2})$ does not contain the leading term $a(Q^{2})^{1}$, and this term has to be subtracted. It turns out that we then get
${\widetilde{B}}[{{\cal D}^{(1)}}](u)_{(p,0)}$ $=$ $\pi\ln\left(1-\frac{u}{p}\right)\;\Rightarrow$ (139a)
${{\cal D}^{(1)}}(Q^{2})_{(p,0)}$ $=$ $\pi\int_{0}^{1}\frac{dt}{t}\frac{t^{p}}{\ln t}\left[a(tQ^{2})-a(Q^{2})\right],$ (139b)
${\cal D}(Q^{2})_{(p,0)}$ $=$ $\pi\int_{0}^{1}\frac{dt}{t}\frac{t^{p}}{\ln t}\left[{\widetilde{a}}_{\nu_{0}}(tQ^{2})-{\widetilde{a}}_{\nu_{0}}(Q^{2})\right].$ (139c)
The corresponding expansion (1) of ${\cal D}(Q^{2})_{(p,0)}$ has $d_{0}(\nu_{0};\kappa)={\widetilde{d}}_{0}(\nu_{0};\kappa)=0$ in such a case (while in the canonical case this coefficient was equal to unity).
As an aside, we mention that quite an analogous proof can be performed for the version of Theorem 4 for the UV renormalon (at u=p<0u=-p<0):
${\widetilde{B}}[{{\cal D}^{(1)}}](u)_{(-p,{\widetilde{s}})}$ $=$ $\frac{\pi}{(p+u)^{{\widetilde{s}}}}$$\Rightarrow$ (140a)
$F_{{{\cal D}^{(1)}}}(t)$ $=$ $\Theta(t-1)\frac{\pi}{\Gamma({\widetilde{s}})\;t^{p}(\ln t)^{1-{\widetilde{s}}}}$ (140b)
where $p>0$ and $0<{\widetilde{s}}<1$. In Ref. renmod , the UV cases for ${\widetilde{s}}=1$ and ${\widetilde{s}}=2$ were investigated, and it turns out that the formula (140b) can be applied even in these cases ${\widetilde{s}}=1,2$.
Appendix C Renormalisation scheme dependence of the expansion coefficients
In this Appendix we present the change of the expansion coefficients $d_{n}(\nu_{0};\kappa)$ that appear in the expansion (1a), or $f_{n}(\nu_{0};\kappa)$ in the expansion (80), for the general case when $\nu_{0}\not=1$ in general. The relations will be equal for the spacelike coefficients $d_{n}$ and timelike coefficients $f_{n}$, because they are based on the same principle of the renormalisation scheme independence of the observables.
The renormalisation scheme dependence of the coefficients and the couplings is parametrised by the parameters $c_{j}\equiv\beta_{j}/\beta_{0}$ ($j=2,3,\ldots$) appearing in the RGE (6). [^33] We note that in the mass-independent schemes, such as $\overline{\rm MS}$, the first two coefficients $\beta_{0}$ and $c_{1}$ are universal. First we note that the running coupling $a(\mu^{2};c_{2},c_{3},\ldots)\equiv a$ has the scheme-dependence given by the following relations (cf. App. A of Stevenson , and App. A of GCRK63 ):
$\frac{\partial a}{\partial c_{2}}$ $=$ $a^{3}+\frac{c_{2}}{3}a^{5}+{\cal O}(a^{6}),$ (141a)
$\frac{\partial a}{\partial c_{3}}$ $=$ $\frac{1}{2}a^{4}-\frac{c_{1}}{6}a^{5}+{\cal O}(a^{6}),$ (141b)
$\frac{\partial a}{\partial c_{4}}$ $=$ $\frac{1}{3}a^{5}+{\cal O}(a^{6}).$ (141c)
When we apply these relations to the physical condition of the scheme dependence of the observable ${\cal D}(Q^{2})$ of Eq. (1a), namely $\partial{\cal D}(Q^{2})/\partial c_{j}$ ($j=2,3,4$), we immediately obtain [^34] Note that $\kappa\equiv\mu^{2}/Q^{2}$ is considered fixed here, and we denote everywhere the values of parameters and coefficients in the $\overline{\rm MS}$ scheme with bar. For simplicity, we also denoted $d_{n}(\kappa;{\bar{c}}_{2},\ldots,{\bar{c}}_{n})$ simply as ${\bar{d}}_{n}(\kappa)$.
$d_{1}(\kappa)$ $=$ ${\bar{d}}_{1}(\kappa),\qquad d_{2}(\kappa;c_{2})={\bar{d}}_{2}(\kappa)-\nu_{0}(c_{2}-{\bar{c}}_{2}),$ (142a)
$d_{3}(\kappa;c_{2},c_{3})$ $=$ ${\bar{d}}_{3}(\kappa)-(\nu_{0}+1)(c_{2}-{\bar{c}}_{2}){\bar{d}}_{1}(\kappa)-\frac{1}{2}\nu_{0}(c_{3}-{\bar{c}}_{3}),$ (142b)
$d_{4}(\kappa;c_{2},c_{3},c_{4})$ $=$ ${\bar{d}}_{4}(\kappa)-(\nu_{0}+2)(c_{2}-{\bar{c}}_{2}){\bar{d}}_{2}(\kappa)-\frac{1}{2}(\nu_{0}+1)(c_{3}-{\bar{c}}_{3}){\bar{d}}_{1}(\kappa)$ (142c)
$-\frac{1}{6}\nu_{0}(c_{2}^{2}-{\bar{c}}_{2}^{2})+\frac{1}{2}\nu_{0}(\nu_{0}+2)(c_{2}-{\bar{c}}_{2})^{2}+\frac{1}{6}\nu_{0}c_{1}(c_{3}-{\bar{c}}_{3})-\frac{1}{3}\nu_{0}(c_{4}-{\bar{c}}_{4}).$
As mentioned, the same relations are valid for the coefficients $f_{n}$. The corresponding scheme transformation relations for the coefficients ${\widetilde{d}}_{n}(\kappa;c_{2},\ldots,c_{n})$ [or ${\widetilde{f}}_{n}(\kappa;c_{2},\ldots,c_{n})$] are obtained then by the direct use of the above transformations and the relations (3) [or: (94)], where we note that the transformation coefficients ${\widetilde{k}}_{n-s}(\nu_{0}+s)$ there are independent of the scheme parameters $c_{j}$ ($j\geq 2$), and even independent of the renormalisation scale parameter $\kappa$, cf. GCAK .
In our implementation of our specific case, we know ${\bar{f}}_{n}$ (at $\kappa=1$) for $n=0,1,2,3$ [cf. Eqs. (81)-(82)], i.e., we implement only the relations (142a)-(142b), for $f_{n}$’s, where $c_{2}$ and $c_{3}$ are the scheme parameters of the MiniMOM scheme with $n_{f}=3$ MiniMOM .